### v1

# library.bib

Abstract
Machine learning methods often struggle with real-world applications in science and engineering due to limited or low-quality training data. In this work, the example of groundwater flow with heat transport is considered; this corresponds to an advection-diffusion process under heterogeneous flow conditions, that is, spatially distributed material parameters and heat sources. Classical numerical simulations are costly and challenging due to high spatio-temporal resolution requirements and large domains. While often computationally more efficient, purely data-driven surrogate models face difficulties, particularly in predicting the advection process, which is highly sensitive to input variations and involves long-range spatial interactions. Therefore, in this work, a Local-Global Convolutional Neural Network (LGCNN) approach is introduced. It combines a lightweight numerical surrogate for the transport process (global) with convolutional neural networks for the groundwater velocity and heat diffusion processes (local). With the LGCNN, a city-wide subsurface temperature field is modeled, involving a heterogeneous groundwater flow field and one hundred groundwater heat pump injection points forming interacting heat plumes over long distances. The model is first systematically analyzed based on random subsurface input fields. Then, the model is trained on a handful of cut-outs from a real-world subsurface map of the Munich region in Germany, and it scales to larger cut-outs without retraining. All datasets, our code, and trained models are published for reproducibility.
1 Introduction
Many real-world systems, across environmental engineering, geoscience, and biomedicine, feature transport phenomena governed by coupled advection–diffusion processes, e.g., pollutant transport through air or water, chemical reactions, or heat transport in flowing media. In such settings, interactions can span large distances, long timescales and many components.
Datasets based on physical measurements are often limited by cost, logistics, or experimental constraints. While numerical simulations can address these limitations to some extent, they become impractically slow or computationally expensive as domain size or resolution increases. Machine learning (ML) models offer fast predictions but typically require large training datasets and often struggle to scale to larger domains. This trade-off underscores the need for data-efficient surrogates that can capture both local and long-range interactions, and scale to arbitrary domain sizes and numbers of interacting components. To this end, we introduce the Local–Global Convolutional Neural Network (LGCNN), a scalable, physics-inspired approach that leverages the structure of advection–diffusion problems by decoupling local and global transport processes.
We apply this approach to the climate change-relevant application of modeling city-wide subsurface temperature fields driven by heat and mass injections of dozens of open-loop groundwater heat pumps (GWHP)[pophillat2020impact, heatpumps, hahnlein2013] in the region of Munich, Germany[geokw22]. A city-wide optimization of the positions of all GWHPs requires calling a prediction model dozens of times with slightly adapted GWHP positions to minimize negative interactions between different heat plumes and pumps. For this scenario, our model should predict the temperature field $T$ in an arbitrarily large domain based on the inputs of the subsurface parameters of a hydraulic pressure gradient $\nabla p$ and a heterogeneous subsurface permeability field $k$, and arbitrary heat pump locations $i$. Labels are generated by expensive simulations, but only for a small number of cases and entirely offline.
As sketched in Figure 1, the LGCNN uses CNNs where they excel, and simple numerical surrogates where CNNs struggle: a first CNN predicts the velocities $\vec{v}$ from local relations of the inputs $p$, $k$ and $i$ (Step 1); a fast numerical component calculates the streamlines $\vec{s}$ based on this (Step 2); a second CNN calculates the temperature distribution $T$ in the whole domain based on all prior inputs and outputs (Step 3). Since Step 2 covers all non-local effects, Step 3 can again take advantage of CNNs to learn local patterns.
Figure 1: Left: Map of $k$ of whole region of Munich. Right: Schematics of our Local-Global CNN-based approach (LGCNN) with 3 physics-inspired steps: CNN ($pki\rightarrow\vec{v}$), simplified solver ($i\vec{v}\rightarrow\vec{s}$), CNN ($pki\vec{v}\vec{s}\rightarrow T$).
We outline related work in Section 2, then define our datasets and metrics in Section 3. In Section 4, we benchmark purely data-driven models and show their performance and limitations on this low-data challenge. To address data scarcity, we introduce LGCNN in Section 5. Section 6 emphasizes generalization and scalability from a single datapoint, while Section 7 demonstrates transfer to real-world subsurface parameters extracted from measurement maps such as in Figure 1.
Contributions
Our method strongly reduces data requirements for modeling complex scenarios, increases reliability, and ensures scalability to larger domains by a domain-specific modular design of physics-aided machine learning for an input-sensitive scenario with long-distance effects. Subsurface temperature prediction serves as our example application, with the approach expected to generalize to related domains.
Limitations
The current approach is limited to two-dimensional and steady-state datasets.
2 Related Work
Deep learning surrogates in scientific modeling can be broadly categorized into purely data-driven architectures, physics-informed or domain-structured models, and highly problem-specific hybrid approaches. In this section, we summarize the developments in each category and existing efforts in GWHP modeling.
Convolutional Neural Networks (CNNs), particularly UNet [CNN15] variants, are widely used for spatial prediction tasks due to their locality-aware architecture of moving-window-like kernels. Applications include flow around airfoils [thuerey_airfoils] and various engineering systems [Jhaveri2022A]. For general reviews on ML in scientific and real-world engineering contexts, see [sharma2021machine, sarker2021machine, angra2017machine].
To improve generalization and reduce data requirements, domain-structured models embed physical priors or symmetries into their architecture. This includes Fourier Neural Operators (FNOs) for efficient multi-scale modeling [fno, choi2024applications], thermodynamics-preserving networks [hernandez2021structure], and rotation-equivariant CNNs (ECNNs) [weiler2023ECNNs]. Physics-Informed Neural Networks (PINNs) go further by replacing data loss with physics-based constraints, allowing for training without labeled data [PINN19, cuomo2022scientific], and have been applied to fluid dynamics and inverse problems [pinn_rao, pinn_sun, cai_physics-informed_2021].
Hybrid approaches combine architectures and integrate domain or physics knowledge to improve general models, for instance, in low-data regimes. Examples include staged CNNs for atmospheric plume dispersion [fernandez2024staged], physics-informed FNOs for traffic flow [thodi2024fourier], CNN–MLP hybrids for optimizing positions of oil wells [yousefzadeh2025determination], and physics-guided CNNs for seismic response prediction [zhang2020physics].
Deep Learning for GWHP Modeling
In the context of groundwater modeling with heat transport, recent work focuses on isolated or pairwise interacting GWHPs. Most use UNet-based architectures, optionally with physics-loss terms [davis2023deep, pelzer2024, Stefania]. Although effective in simple, homogeneous aquifers, these models rely on large training datasets or simulated inputs and have not been scaled to city-wide domains with many pumps. UNets are the dominant architecture because of their strong performance on spatial data, making them the baseline for comparison.
We briefly evaluated PINNs and FNOs as alternatives. PINNs, while promising for low-data fluid dynamics tasks [pinn_rao, pinn_sun, cai_physics-informed_2021, pdebench], struggled with complex scenarios (cf. [krishnapriyan2021characterizing]) such as heterogeneous media and discontinuous source terms in preliminary tests. FNOs, though capable of modeling local and global dependencies, struggled with large domain sizes due to high memory requirements. Additionally, they performed poorly in the presence of multiple sources, consistent with [liu2024neural].
3 Datasets and Metrics
Datasets
Inputs for the neural networks consist of a heterogeneous permeability field $k$, an initial hydraulic pressure field $p$, and a one-hot-encoded field of heat pump positions $i$. The (interim) labels of velocity $\vec{v}$ and temperature $T$ fields are simulated, in our case with Pflotran [pflotran], until a quasi-steady state is reached after $\approx$27.5 years simulated time [bawue_arbeitshilfe]. All data are normalized to $[0,1]$, and stored in PyTorch format for training.
We generate two types of datasets, one is based on synthetic, the other on real permeability fields $k$. Both cover a $12.8\times 12.8$ km2 domain with 2 560×2 $\times 2\,560$ 2 560 cells. The baseline dataset uses Perlin noise [perlin] to generate random, heterogeneous $k$. Three simulations (3dp) with different fields for $i$ and $k$ are run, generating one datapoint each for training, validation, and testing, plus a simulation 4$\times$ larger to assess scalability. For training vanilla neural networks in Section 4, we generate an additional dataset of 101 datapoints (101dp), split into 73:18:10 for train:val:test. Runtimes are $\approx$27 hours per simulation and 123 hours for the larger domain.
The more realistic dataset builds on $k$ fields cut from maps of borehole measurements in the Munich region[lfu_gepo]. This dataset consists of four simulations (three for training and one for validation) and one larger simulation for testing scalability. Due to constraints of the available measurements, see Figure 1, the large-scale simulation extends only in length, resulting in a rectangular domain twice the length. Runtimes range from 38 to 91 hours (average 58 hours), and 134 hours for the scaling-test domain. Variation across the dataset stems from different heat pump placements in $i$ and from the specific regions where $k$ is extract from the available measurement domain.
For detailed information on simulation procedures and hydro-geological parameters, see Section A.1; for hardware specifications see Section A.4.
Performance Metrics
We evaluate model accuracy per output dimension ($v_{x},v_{y},T$) separately using Mean Absolute Error (MAE) [naser2023error], Mean Squared Error (MSE) [naser2023error], Maximum Absolute Error ($L_{\infty}$) Structural Similarity Index Measure (SSIM) [ssim], and application-driven metrics of Percentage Above Threshold (PAT) and Visual Assessment. PAT measures the percentage of cells where the absolute error of the predicted temperature exceeds the threshold of $0.1$C, corresponding to measurement precision [kabellichtlot], and is only applicable to temperature predictions. All metrics other than SSIM are applied after re-normalization to the original data ranges to obtain physically meaningful results in [°C] or [meters/year].
4 Application of Purely Data-Driven Approaches
In this section, we present the performance of two purely data-driven approaches for predicting $T$ from inputs of $p,k,i$: a UNet [CNN15] and the domain-decomposition-based UNet variant, called DDUNet [DDUNet].
Methods
The fully convolutional vanilla UNet has an encoder-decoder architecture with skip connections between layers of equal spatial dimensions. Convolutional networks excel in extracting local dependencies from spatial data with locally repeating patterns. Hence, in theory, one datapoint covering a large domain could be enough to generalize, if only local dependencies occur. Global patterns would require a large receptive field and, hence, more data. The DDUNet is designed to handle large-scale domains under GPU memory constraints. It combines a UNet architecture with domain decomposition principles by partitioning the spatial domain into subdomains that are distributed across multiple GPUs. We use a decomposition into $2\times 2$ subdomains. [^1] Conceptually, DDUNet reduces to a standard UNet when trained with a single subdomain. Global context is communicated between subdomains via coarse feature maps through a learned communication network, maintaining spatial consistency and providing a large receptive field while reducing memory demands. The hyperparameters of the models are optimized according to Section A.2.1.
Experiments
Experiments for the vanilla UNet are conducted on the baseline datasets with 3dp and 101dp, those with DDUNet directly on the 101dp dataset, see Section A.3. Both models are applied to the test of 3dp and to the scaling datapoint, for direct comparability to LGCNN. UNet exhibits poor performance when trained on the limited 3dp dataset (UNet3dp), as demonstrated in both Figure 4 and Table 1. Similar performance would be expected for the DDUNet when applied to this sparse dataset. When trained on 101dp, however, both demonstrate robust predictive and scaling capabilities. The advantage of DDUNet over vanilla UNet is a reduction of memory usage and runtime, see section A.3.
[…]
Table 1: Performance metrics for predicting $T$ from $pki$ with UNet and DDUNet. Errors in [°C], MSE in [°C2], PAT in [%], and SSIM unitless. Inferred on the test of 3dp and the scaling datapoint.
Model Test Scaling
L MAE MSE PAT SSIM L MAE MSE PAT SSIM
UNet3dp 4.8642 0.1314 0.0492 39.05 0.5794
UNet101dp 4.3985 0.0473 0.0100 13.63 0.9827 4.3426 0.0202 0.0033 4.33 0.9955
$2\times 2$-DDUNet101dp 3.4257 0.0548 0.0128 17.42 0.9804 4.1806 0.0235 0.0052 6.11 0.9940
5 Methodology
In data-scarce settings, performance can improve when the structure of a model reflects the underlying physics, cf. Section 2. This section hence outlines the relevant physical processes, motivates our LGCNN method, and discusses a few details of LGCNN’s individual components.
Physics of Groundwater Flow with Heat Transport
Transport of heat induced by heat injections of GWHPs in the subsurface is an advection–diffusion process. Since the Péclet number, i.e., the advection-diffusion ratio, is high in our application (Pemuch-$\gg$1, more in Section A.1), the system is advection-dominated. Advection is largely governed by the global hydraulic pressure gradient, driving flow from higher to lower regions. Locally, the flow paths are influenced by the spatial distribution of permeability $k$, cf. Figure 2. As a result, small changes in $k$ can lead to significant differences in flow paths further downstream, demonstrating high input sensitivity.
One-Way Coupled Approach: Local-Global CNN (LGCNN)
In a fully coupled physical model, flow field computation and heat transport along streamlines [streamline] starting at heat pumps (advection) and their widening (diffusion) must be solved monolithically. By splitting effects into local and global and working with steady-state simulated $\vec{v}$ during training, we can reduce the coupling to one direction without significant loss in accuracy. Our simplified physical pipeline consists of three steps: (1) compute a steady-state flow field $\vec{v}$ from initial subsurface parameters $p$ and $k$, with $i$ encoding mass influx around GWHPs; (2) transport injected heat along streamlines governed by $\vec{v}$ until quasi-steady state[bawue_arbeitshilfe]; and (3) apply plume widening to these heat paths, informed by soil diffusivity and $\vec{v}$, to approximate diffusion effects. The resulting one-way coupled LGCNN approach (see Figure 1) can formally be described as:
  1. Step 1
    Velocities (local) We employ a CNN to predict the velocity field $\vec{v}=\left(v_{x},v_{y}\right)$ from $p$, $k$, and $i$:
    $\text{CNN}\left(p,k,i\right)=\vec{v}.$ (1)
  2. Step 2
    Streamlines (global) Based on the predicted velocities $\vec{v}$ from Step 1, we compute streamlines $\vec{s}$ originating from all pump locations in $i$ with an initial value problem (IVP) solver:
    $\text{IVP}\left(i,\vec{v}\right)=\vec{s}.$ (2)
  3. Step 3
    Temperature (local) A second CNN predicts the temperature field $T$ from the initial inputs $p$, $k$, $i$, the predicted velocities $\vec{v}$ of Step 1 and streamlines $\vec{s}$ calculated in Step 2:
    $\text{CNN}\left(p,k,i,\vec{v},\vec{s}\right)=T.$ (3)
The model outputs the steady-state temperature field $T$ and, using the same inputs $p$, $k$, and $i$ as a simulation, serves as its direct surrogate.
CNN Models in the Local Steps
The local Steps 1 and 3 are approximated by a UNet[CNN15, pelzer2024], see Section A.2 for details of our architecture. It is beneficial to omit zero-padding to enforce shift invariance and reliance on purely local information; this approach is also validated by Islam et al. in [Islam2020How]. Due to the input sensitivity of the entire problem, each step is trained independently on simulated input data. In particular, the CNN in Step 3 is trained on the simulated velocities $\vec{v}_{\text{sim}}$ and streamlines based on these fields. Only during inference, all steps are applied sequentially, i.e., Step 3 uses the outputs of Steps 1 and 2, including the predicted $\vec{v}_{\text{pred}}$ as inputs.
The training is accelerated by partitioning the data into overlapping patches, which increases the effective number of datapoints while reducing their spatial size. Combined with the streamlines’ embedding and localizing of global flow patterns for the fully convolutional (locally) acting CNN, this training data enrichment allows the model to train effectively on very little data, sometimes as little as a single simulation run, while generalizing to unseen and even to larger domains. Validation and testing are conducted on independent simulation runs, with the full domain being processed as a whole instead of patches during inference.
The architecture, input choices, and training hyperparameters — including the patch size and overlap in the datapoint partitioning — were optimized through a combination of manual and automated tree-based search using Optuna [optuna_2019]; see Section A.2.1 for details.
Streamline Calculation and 2D-Embedding in the Global Step
Streamlines are calculated by solving the initial value problem (IVP)
$\frac{dy}{dt}=\vec{v}(y),\text{ with }y(t_{0})=y_{0},$ (4)
with a lightweight numerical solver for each heat pump, where $y_{0}$ represents the location of a heat pump in $i$. We employ the solve_ivp function from SciPy [scipy] with an implicit fifth-order Runge-Kutta method [scipyRadaux2014] and 10 0001000010\,00010 000 time steps of one day, corresponding to the simulation time of forming a heat plume.
The computed streamlines $y$ represent a sequence of positions in 2D, that a particle - injected at a heat pump’s location - would traverse based on $\vec{v}$. The corresponding grid cells are assigned values that fade linearly from 1 to 0, reflecting the time it takes to reach each position, similar to a soft one-hot encoding.
To address the sensitivity of this process with respect to $k$ and $\vec{v}$, we calculate a set of outer streamlines $s_{o}$. For this, we perturb each heat pump’s location $y_{0}$ by 10 cells orthogonal to the global flow direction, i.e., $\nabla p$, and compute the corresponding streamlines. By incorporating $\vec{s}=\left(s,s_{o}\right)$, the subsequent Step 3 obtains a spatial, localized representation of flow paths. In future work, we aim to explore probabilistic perturbations to compute the mean and standard deviation of streamlines, which, while being computationally more expensive, could be efficiently parallelized on GPUs.
6 Results on Synthetic Permeability Fields
This section evaluates our model’s performance on the baseline dataset 3dp and on a scaling test datapoint; cf. Section 3. In Section 6.1, we present metrics for isolated velocity (Step 1) and temperature (Step 3) predictions, and for the full pipeline, demonstrating the potential of our approach. In ablation studies (Section 6.2), we motivate model design choices. Additional metrics and visualizations can be found in Section A.3.
Table 2: Metrics for predicting $\vec{v}_{\text{pred}}$ in Step 1 and in ablation studies of Step 1 on test and on scaling data from our baseline 3dp dataset with random $k$. In [m/y], except for MSE in [m2/y2], SSIM unitless.
Output Case L MAE MSE SSIM
LGCNN Step 1
$v_{x}$ test 190.8046 22.3178 972.5668 0.9911
$v_{y}$ test 256.2519 32.7444 2031.3357 0.9812
$v_{x}$ scaling 294.0457 24.9261 1204.1154 0.9911
$v_{y}$ scaling 367.6891 26.2795 1463.8218 0.9820
Ablation study Train without data partitioning
$v_{x}$ test 202.6397 23.8584 1144.4479 0.9913
$v_{y}$ test 332.9557 35.9249 2490.4104 0.9789
[…]
$v_{x}$ test 1047.8464 12.7036 896.2758 0.9016
$v_{y}$ test 2594.9504 14.3878 2317.6863 0.9903
$v_{x}$ scaling 659.8439 5.2152 79.2508 0.9699
$v_{y}$ scaling 456.1904 5.7492 83.9735 0.9948
[…]
$v_{x}$ test 2453.5881 13.6735 1367.9061 0.9325
$v_{y}$ test 2029.0269 15.1332 2221.4226 0.9919
$v_{x}$ scaling 421.4186 8.1184 125.9511 0.9314
$v_{y}$ scaling 400.2490 7.4200 135.7929 0.9955
Table 3: Metrics for predicting $T$ in Step 3, in the full pipeline, and in ablations studies of Step 3 on test and on scaling data from our baseline 3dp dataset with random $k$. In [°C], except for MSE in [°C2], PAT in [%], SSIM unitless.
Case L MAE MSE PAT SSIM
LGCNN Step 3, i.e., test on $\vec{v}_{\text{sim}}$
test 2.8990 0.0347 0.0041 7.54 0.9304
scaling 3.0250 0.0168 0.0014 2.05 0.9510
Pipeline, i.e., test on $\vec{v}_{\text{pred}}$
test 4.2120 0.0905 0.0307 28.92 0.7637
scaling 4.9366 0.0413 0.0141 10.87 0.8654
Ablation study Train in sequence
test 4.0804 0.0901 0.0289 29.44 0.7553
Include zero-padding
test 2.3391 0.0418 0.0048 9.19 0.8838
Train without data partitioning
test 3.6420 0.0351 0.0043 7.41 0.9245
[…]
test 4.4658 0.0416 0.0067 11.69 0.9835
scaling 4.3489 0.0176 0.0022 3.07 0.9954
[…]
test 4.2772 0.0376 0.0058 10.96 0.9915
scaling 3.9946 0.0192 0.0025 4.73 0.9968
6.1 Performance of LGCNN
Isolated Steps 1 and 3
Performance metrics for Step 1 are shown in Table 3, for Step 3 (trained and tested on simulated $\vec{v}_{\text{sim}}$) in Table 3. The combination of quantitative (MAE of (22.3, 32.7) m/y and 0.03 °C) and qualitative results through a visual assessment of Figure 3 demonstrate a strong performance with heat plumes forming according to $k$.
Figure 3: $pki\vec{s}\vec{v}_{\text{sim}}\rightarrow T$ (baseline dataset (3dp), test). 1st column: Label $T$. 2nd-3rd column: LGCNN Step 3, trained without and with zero-padding.
Full Pipeline
Testing the full pipeline, i.e., Steps 2 and 3 are based on $\vec{v}_{\text{pred}}$, results in a higher test MAE of 0.09 °C compared to an isolated Step 3 (see Table 3). This was expected as the isolated Step 3 uses the true, simulated velocity fields $\vec{v}_{\text{sim}}$ as inputs. Visual assessment of representative predictions in Figure 2 reveals physically plausible heat plumes in terms of shape, extent, and heat magnitude. Deviations in streamlines arise from smaller errors in the velocity predictions from Step 1 $\vec{v}_{\text{pred}}$, highlighting the input sensitivity, especially near bifurcations, where small perturbations in $\vec{v}$ can lead to an alternative path around a clay lens in $k$. Compared to fully data-driven models trained on 101dp (see Table 1), our model achieves a similar accuracy while training on only 1 versus 73 training datapoints, which strongly reduces computation time (see Table 16 for details).
Scaling Test
We can already expect that our model is able to scale to larger domains, as we train only on local patches of the spatial domain and achieve high test accuracies on the full domain during validation and testing; see also Section 6.2 Training Data Partitioning for more details. To further demonstrate the scaling, we test our model on a domain of 4$\times$ the size of the training domain with the same number of heat pumps, hence with a lower density of GWHPs/km2; cf. Figure 4.
[…]
We obtain a comparable MAE of 0.02 °C for Step 3, and an MAE of 0.04 °C for the full pipeline, see Table 3. A visual assessment of Figure 4 yields similar qualitative behavior as the previous results in Figure 2.
6.2 Ablation Study
Training in Sequence
[…]
Zero-Padding
Removing zero-padding from all convolutional layers yields smaller but cleaner output fields, as shown in Figure 3. As a result, we obtain improved performance, reflected in lower PAT and higher SSIM scores in Table 3.
Partitioning Training Data
Hyperparameter search shows the best performance for partitioning the domain during training into 20 736 overlapping patches of 256$\times$256 cells for Step 1 and 82 944 patches for Step 3. Although this increases per-epoch compute time compared to full-domain training, it significantly reduces the number of epochs needed to reach comparable test performance (cf. Tables 3 and 3), resulting in an overall training time reduction of 67–90% (cf. Table 15).
[…]
We individually replace each of the three steps with an optimized (cf. Section A.2.1) vanilla approach, i.e., UNet or DDUNet trained on 101dp. For Step 2, both fail completely; results for Steps 1 and 3 are summarized in Tables 3 and 3. In Step 1, UNet slightly outperforms DDUNet, while the reverse holds for Step 3. A visual comparison of the scaling datapoint in Figure 4 reveals that increasing training data from 3dp to 101dp has minimal impact on the performance in Step 3.
Modifying Streamline Inputs of Step 3
We evaluate the influence of Step 3’s inputs, focusing on the effect of modifying or omitting individual components of the streamlines $\vec{s}=\left(s,s_{o}\right)$. When both the central streamlines $s$ and offset streamlines $s_{o}$ are excluded, predicted plumes fail to follow flow paths beyond the CNN’s receptive field. Omitting only $s_{o}$ results in overly narrow plumes, whereas excluding only $s$ produces overly smeared temperature fields. Including both but removing the fading of values along the streamlines leads to diffuse background temperatures and less localized plumes. Overall, the Huber validation loss increases by 32–132%, depending on the experiment. For visual, quantitative, and extended results, we refer to Section A.3.
7 Domain Transfer to Real Permeability Fields
Finally, we investigate whether our approach transfers to the realistic dataset in Section 3, which is based on real-world permeability fields $k$ that are cut from geological maps of Munich. Since these maps cover a limited area, the number of potential datapoints is very small. We compare performance for temperature prediction of Step 3 and the full pipeline on the validation and scaling datapoints. Results are presented in Tables 5 and 5; additional metrics are in Section A.3.
Table 4: Metrics for predicting $\vec{v}_{\text{pred}}$ in Step 1 on validation and scaling data from our realistic dataset. In [m/y], except for MSE in [m2/y2], SSIM unitless.
Output Case L MAE MSE SSIM
Step 1
$v_{x}$ val 106.8607 15.4095 380.3762 0.9939
$v_{y}$ val 74.5570 10.6605 148.2475 0.9993
$v_{x}$ scaling 459.1620 110.0078 13570.0000 0.9462
$v_{y}$ scaling 240.9406 17.6051 616.5186 0.9965
Table 5: Metrics for predicting $T$ in Step 3 and in the full pipeline on validation and scaling data from our realistic dataset. In [°C], except for MSE in [°C2], PAT in [%], SSIM unitless.
Case L MAE MSE PAT SSIM
Step 3, i.e., test on $\vec{v}_{\text{sim}}$
val 0.8222 0.0175 1.01e-3 2.18 0.9672
scaling 0.8052 0.0189 8.21e-4 0.92 0.9497
Full pipeline, i.e., test on $\vec{v}_{\text{pred}}$
val 2.3194 0.0841 2.75e-2 27.79 0.7510
scaling 2.0511 0.0394 4.42e-3 10.02 0.8708
Methodological Adaptations
Slight adjustments to chunk size, overlap, data split, and architecture yield better performance compared to those used on the baseline dataset, see Section A.2.1 for details. Visual inspection of input $k$ in Figure 5 reveals fewer but larger-scale features. Therefore, training can benefit from a larger spatial context, which requires more training data. We increase the number of training samples by using three out of the four available datapoints for training and one for validation. For the streamlines computation, we switch to an explicit fourth-order Runge–Kutta scheme, which proved stable and sufficient for training and inference due to the lower frequency in $k$.
Performance and Scaling Tests
In Step 3, we reach an even lower maximum absolute error $L_{\infty}$ of 0.82 °C compared to 2.90 °C of our initial model on the baseline dataset (Section 6). Other metrics confirm a similar performance; see Table 3. In the scaling test, most errors are comparable too. For the full pipeline, errors are slightly higher on the validation and scaling data compared to an isolated Step 3, consistent with our observations on the baseline dataset (Table 3). Physical consistency, which is our primary objective, remains strong: Shape, magnitude, and connectivity of the predicted heat plumes are well preserved, even in the scaling test (Figure 5). Local deviations are primarily due to differences between $\vec{v}_{\text{sim}}$ and $\vec{v}_{\text{pred}}$, as evident from the discrepancy between Step 3 and full pipeline outputs.
Figure 5: $pki\rightarrow T$ (realistic dataset, scaling). 1st & 2nd column: Label $T$ and input $k$. 3rd-4th column: LGCNN (isolated Step 3 and full pipeline).
8 Conclusions and Future Work
In real-world scientific applications, data is often very limited. At the same time, models are expected to generalize and scale to large domain sizes, pushing standard data-driven models to their limits. To address this issue, we proposed the LGCNN model, a hybrid, physics-inspired CNN approach. In the prediction of groundwater temperature fields under arbitrary heat pump configurations and real permeability data, we identify long-range interactions driven by advection as the problematic process for classical deep learning models, such as vanilla CNNs. We replace this part of the prediction with a simple numerical solver, resulting in a sequence (CNN 1 – numerical solver – CNN 2). While this setup is specific to groundwater heat transport, the general approach should be applied to a wider range of applications. Finally, we stress that traditional error metrics can obscure effects like input sensitivity; therefore, we additionally rely on visual assessments and problem-specific metrics such as PAT. In future work, we will address dataset limitations, investigate input sensitivity, and accelerate streamline computation.
Acknowledgements
We thank the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) for supporting this work by funding - EXC2075 – 390740016 under Germany’s Excellence Strategy. We acknowledge the support by the Stuttgart Center for Simulation Science (SimTech). We also acknowledge the use of computational resources of the DelftBlue supercomputer, provided by Delft High Performance Computing Centre [DHPC2024]. We thank Kai Zosseder and Fabian Böttcher at the Technical University of Munich for their application-related input.
[…]
Data and Code Accessibility
Our raw data and pretrained models are publicly available on DaRUS https://darus.uni-stuttgart.de/dataverse/ipvs_lgcnn/ or more specifically:
Raw datasets:
Pretrained models (including hyperparameters):
[…]
  • LGCNN on real permeability fields, 4dp: https://doi.org/10.18419/darus-5082
Our code for data generation (https://github.com/JuliaPelzer/Dataset-generation-with-Pflotran) and model training (including dataset preparation and evaluation routines) (for LGCNN and data preparation: https://github.com/JuliaPelzer/Heat-Plume-Prediction/tree/AllIn1/LGCNN/release25, for DDUNet: https://github.com/corne00/DDUNetForHeatplumePrediction) is on GitHub. Follow the instructions in the respective readme files to prepare the data and train or infer a model.
Appendix A Appendix
A.1 Simulation Setup and Transport Regime
In Section 3, we describe how the two datasets are generated using the subsurface simulation software Pflotran [pflotran], which solves the coupled mass and energy conservation equations. Here, we provide additional technical details and modeling assumptions relevant for reproducibility. Furthermore, we show that for the selected parameters, the heat transport in our system is dominated by advection by a theoretical analysis.
Hydro-geological, Operational and Simulation Parameters
Our data was generated on 12.8km×$\times$12.8km×$\times$5m with a cell size of 5m×$\times$5m×$\times$5m for the baseline simulations. To test scalability, we also simulate a domain that is twice as large in both $x$- and $y$-dimension for the baseline dataset, but only twice as large in $y$-dimension for the more realistic dataset due to dimension restrictions of the available input data.
The baseline dataset uses a constant hydraulic pressure gradient $\nabla p$ of 0.003 [pelzer2024, geologica]; permeability field $k$ is generated using Perlin noise [perlin] within (1.02E-11, 5.10E-9) m2, and $100$ randomly placed heat pumps, which are all operating with a 5 i.e., C injection temperature difference compared to the surrounding groundwater and an injection rate of 0.00024 m3/s. All values are based on realistic parameter ranges in the region of Munich[geologica, geokw22].
For the realistic dataset, instead of random permeability fields $k$, we use permeability field data that is derived from borehole measurements in the Munich region[lfu_gepo]. Furthermore, based on subsurface measurements, we set $\nabla p=0.0025$ for these simulations. All other simulation parameters remain identical to the baseline. Other subsurface and operational parameters are taken directly from [pelzer2024], which also includes additional information about solver setup and boundaries. For mathematical details of the governing equations, we refer the reader to [pflotran_theory, anderson2007introducing, delleur2016elementary].
Simplifications
For simplifications, we refer to the (hydraulic) pressure field $p$ throughout the paper. In our paper, this field is the initial pressure field defined through the measured hydraulic head and is not the true pressure field at quasi-steady state with spatial details resulting from permeability $k$ variations in the domain and mass injections at the heat pump positions. The true pressure field is only available after simulation (which we are avoiding for our approach) and strongly interacts with the velocity field.
Péclet Number
To quantify whether heat transport in our system is dominated by advection or diffusion, we compute the dimensionless Péclet number [rapp2017peclet], which is defined as
$\text{Pe}=\frac{L\cdot v}{\alpha},$
with the characteristic length $L$, the local velocity $v$, and the thermal diffusivity $\alpha$, defined as
$\alpha=\frac{k}{\rho c_{p}},$
with $k$ the thermal conductivity, $\rho$ the density, and $c_{p}$ the specific heat capacity. We take the parameters of the solid phase of our simulation inputs of
  • Thermal conductivity: $k=0.65\,\mathrm{W/(m\cdot K)}$
  • Density: $\rho=2800\,\mathrm{kg/m^{3}}$
  • Specific heat capacity: $c_{p}=2000\,\mathrm{J/(kg\cdot K)}$
This yields a thermal diffusivity of:
$\alpha=\frac{0.65}{2800\cdot 2000}\approx 1.16\times 10^{-7}\,\mathrm{m^{2}/s}$
The velocity values are derived by simulations, taken in the direction of flow (longitudinal) from the realistic $k$-dataset:
  • Maximum: $1200\,\mathrm{m/year}\approx 3.8\times 10^{-5}\,\mathrm{m/s}$
  • Minimum: $44\,\mathrm{m/year}\approx 1.4\times 10^{-6}\,\mathrm{m/s}$
  • Mean: $330\,\mathrm{m/year}\approx 1.04\times 10^{-5}\,\mathrm{m/s}$
[…]
$Pe=\frac{L\cdot v}{\alpha}\approx L\cdot 10^{2}\ \approx 10^{5}$
The interpretation of the Peclet number is given by the following:
  • $Pe\ll 1$: Diffusion dominates
  • $Pe\approx 1$: Diffusion and advection similarly dominate
  • $Pe\gg 1$: Advection dominates
Therefore, $\text{Pe}\approx 10^{5}$ indicates that in the simulations, the heat transport is advection-dominated at larger scales.
A.2 Neural Networks, Hyperparameters and Training Details
This section provides implementation and training details for all neural network models used in this work. We begin with a description of the baseline UNet architecture, which also forms the foundation of both LGCNN and DDU-Net. Then, we outline the hyperparameter optimization process and present the final configurations per model and dataset. All models are trained using PyTorch[paszke2017automatic].
UNet Architecture
The UNet architecture used in this work is based on the original design presented by Ronneberger et al. [ronneberger2015u], with several modifications to tailor it to the specific needs of our task of predicting temperature fields. In this section, we introduce the key concepts that define the architecture and explain how they contribute to the model’s design. These concepts will be employed in the hyperparameter search to identify the most suitable configuration, taking into account memory and data limitations.
The UNet architecture essentially consists of a series of UNet blocks. Each block consists of the layers of "Convolution - Activation - Convolution - Norm - Activation - Convolution - Activation" with convolutional layers, a batch normalization layer, and activation functions. After each block, either an upsampling or a downsampling operation is applied via "Max Pooling 2D" or "Transposed Convolution 2D" with stride 2. The depth of the UNet refers to the number of UNet blocks in both the encoder and decoder. The number of initial features refers to the number of feature maps generated by the first downsampling block. Each downsampling block produces twice as many output feature maps as input feature maps, while each upsampling block reduces the number of feature maps by half. The number of convolutions per block denotes how many convolutional layers are applied within each block, while kernel size specifies the size of the kernels used in the convolutional operations. Additionally, we explore different activation functions (ReLU, tanh, sigmoid, and LeakyReLU) and various normalization strategies (batch normalization, group normalization, and no normalization). The UNet block exists in 2 variants, depending on the hyperparameter repeat inner: If repeat inner = False, the block looks like this "Convolution - Norm - Activation", if it is True, the block looks as described above.
The training process also involves several hyperparameters. The learning rate controls how quickly the model adjusts its parameters during training. The weight decay parameter helps prevent overfitting by penalizing large weights in the model. Furthermore, the Adam optimizer [kingma2014adam] is employed as optimizer. For the realistic permeability field, we additionally introduce the optimizable hyperparameter optimizer switch, which, when enabled, switches the optimizer to LBFGS after 90 epochs.
During inference, each datapoint is processed as a whole, but during training, they are loaded in smaller patches cut out from the datapoint. Optimized hyperparameters include their overlap, i.e., inverse of skip per direction, and their size, called box length. The data split is untouched by this, i.e. one simulation run per train / val / test separately. This is important to not overlap test patches with training ones.
A.2.1 Hyperparameters and Hyperparameter Optimization
We optimize the hyperparameters of our architecture, training process and data loading in several rounds with Optuna [optuna_2019] and additional manual tweaking. Optuna performs optimization using a tree-structured Parzen estimator algorithm. An overlook of all considered hyperparameters, their ranges and our final choice, as well as the hyperparameters fixed during optimization is provided per used model architecture and dataset, e.g., purely data-driven approaches, LGCNN on individual steps or the full pipeline, on datasets of random $k$ versus realistic $k$.
Vanilla UNet
[…]
The inputs to the model are $pki$ (pressure field, permeability field, and location of heat pumps). For training, the Mean Absolute Error (MAE) loss function was used, and optimized with the Adam optimizer. The learning rate is set to $1\times 10^{-5}$. The model is trained for 10,000 epochs.
Hyperparameters for purely data-driven approaches
The values considered during hyperparameter optimization for the UNet and DDUNet, trained on 73 datapoints and performed using Optuna [optuna_2019], along with the best settings found, are listed in Table 6. Certain hyperparameters were fixed: we set the number of epochs to 750, with an early stopping criterion based on validation loss and a patience of 80 epochs. Additionally, we note that some hyperparameter combinations (e.g., 32 initial features, depth 6, and 3 convolutions per layer with a kernel size of 7) caused memory issues, leading to their exclusion from the hyperparameter search.
Table 6: Overview of used hyperparameters for the UNet101dp and ${2\times 2}$ DDUNet101dp their search ranges (if applicable), and best values across training stages. Note that the number of communicated feature maps in the vanilla UNet is simply an extra convolution layer in the coarsest part of the UNet (without communication).
Hyperparameter Range $pki\rightarrow v_{x}v_{y}$    (Step 1) $kiv_{x}v_{y}ss_{o}\rightarrow T$    (Step 3) $pki\rightarrow T$    (Full)
UNet101dp ${2\times 2}$DDUNet101dp UNet101dp ${2\times 2}$DDUNet101dp UNet101dp ${2\times 2}$DDUNet101dp
Dataset
Batch size (train) 4, 6, 8 4 6 6 4 6 6
Include pressure field True, False False False - - False False
Encoder-decoder properties
Depth 4, 5, 6 6 5 6 6 5 5
No. initial features 8, 16, 32 8 16 8 8 8 16
No. convs. per block 1, 2, 3 1 1 3 3 3 3
Kernel size 3, 5, 7 7 7 3 5 7 7
Communication Network
No. comm. feature maps 64, 128, 256 64 128 256 64 256 256
Training
Learning rate [1e-5, 1e-3] 0.00024 0.00100 0.00017 0.00030 0.00024 0.00024
Weight decay 0.0, 0.001 0.0 0.0 0.0 0.0 0.0 0.0
Train loss MSE, L1 MSE MSE L1 MSE MSE MSE
After the hyperparameter search, the values corresponding to the best-performing model (based on Huber loss for the validation dataset) were selected. With these values fixed, five models were trained using different randomly sampled initializations to evaluate sensitivity to random initialization, for these values see Tables 11 and 10.
LGCNN Hyperparameters - on Random Permeability
The values considered during hyperparameter optimization with Optuna [optuna_2019] and the best settings found for both steps of LGCNN are listed in Table 7. Although the optimization was originally run for 100 epochs, the optimum was consistently found within the first 25 epochs. Therefore, to reduce computation cost, we therefore conservatively lowered the maximum number of epochs to 50. This adjustment does not affect any of the reported metrics in the paper.
Table 7: LGCNN-Random $k$: Hyperparameter optimization: Parameter ranges and best configurations.
Parameter Range 1st step 3rd step
Parameters of the Dataset
inputs $v:$ $(p,i,k)$ $pik$ $iv_{x}v_{y}ss_{o}k$
$T:$ $(i,v_{x},v_{y},s,s_{o},k)$
skip per direction $v:$ 4, 8, 16, 32, 64 16 8
$T:$ 8, 16, 32, 64
box length 64, 128, 256, 512 256 256
Parameters of Training
loss function (training) MAE, MSE MSE MAE
optimizer Adam, SGD Adam Adam
Parameters of the Network
No. initial features 8, 16, 32, 64, 128 32 32
kernel size 3, 4, 5 5 4
depth $v:$ 1, 2, 3, 4 4 4
$T:$ 1, 2, 3
Fixed parameters for this hyperparameter search are the learning rate (fixed at $10^{-4}$), ReLU as activation function, the batch size of 20, and the use of a batch normalization layer within the inner blocks of the UNet architeture. The validation loss used for selecting the optimal model is the MAE.
LGCNN Hyperparameters - on Real Permeability
The values considered during hyperparameter optimization on the dataset with a more realistic permeability field were selected using Optuna [optuna_2019], and are summarized in  Table 8, along with the best configurations found for both steps of the LGCNN. The optimization was run for up to 100 epochs. For more background on the network architecture and the various hyperparameters, cf. A.2. Several hyperparameters were fixed during this process. These include a constant learning rate schedule, an Adam optimizer with a weight decay of $10^{-4}$, and, when enabled, a switch to LBFGS after 90 epochs. Fixed architectural parameters include a convolutional stride and dilation of 1. During training, the inputs were cut out from the full datapoints. For model comparison, the validation loss was consistently computed using the Huber loss.
Table 8: LGCNN-Real $k$: Hyperparameter optimization: Parameter ranges and best configurations.
Parameter Range model $\vec{v}$ model $T$
Parameters of the Dataset
inputs $v:$ $ki,pik,gik,gk,pk$ $pik$
$T:$ $iv_{x}v_{y}ss_{o}k$ $iv_{x}v_{y}ss_{o}k$
batch size 2, 4, 8, 16 8 8
skip per direction 256, 128, 64, 32, 16, 8 8 8
box length 1280, 640 1280 1280
Parameters of Training
loss function (training) MSE, MAE MSE MSE
optimizer switch True, False False False
learning rate 1e-3, 5e-4, 1e-4, 5e-5 1e-3 1e-4
Parameters of the Network
No. initial features 8, 16, 32 32 16
kernel size 3, 5 3 5
depth 4, 5, 6 5 6
repeat inner True, False False False
activation function relu, tanh, sigmoid, leakyrelu relu relu
layer norm batch-, group-, None batch- batch-
A.3 Additional Experimental Results
This section provides additional experimental results. While the main results section focused only on the test and scaling datasets, we also include here the metric values on the training and validation datasets. For completeness and easier comparison, the test and scaling metrics are re-listed as well.
Purely data-driven UNet, DDU-Net: Metrics of training and ablation study
Table 9 presents all the metrics for predicting the temperature field directly from the inputs $pki$ using a data-driven approach, evaluated on the training, validation, and test datasets. The results are provided for several models: (1) UNet trained on only 1 datapoint and tested and validated on 2 additional datapoints, (2) a UNet trained on 73 datapoints (73-18-10 train-validation-test split), and (3) a DDUNet trained on the same 73 datapoints dataset, operating on $2\times 2$ subdomains. Furthermore, to assess the model’s sensitivity to random initialization, the training of the same architecture was repeated five times for the most relevant models. Based on these repetitions, the mean and standard deviation of the performance metrics were computed using the following equations:
$\bar{x}=\frac{1}{n}\sum_{i=1}^{n}x_{i}\quad\text{and}\quad\sigma=\sqrt{\frac{1}{n-1}\sum_{i=1}^{n}(x_{i}-\bar{x})^{2}}$
where $x_{i}$ denotes the metric value from the $i$-th training run, and $n=5$ is the number of runs. These results are summarized in Table 10 and Table 11. The choice of $n=5$ was made empirically to balance computational effort and statistical reliability. The standard deviations in  Tables 11 and 10 were used as validation: they are neither excessively large (indicating instability) nor unrealistically small (indicating insufficient sampling).
In addition to testing the UNet and DDUNet trained on 73 datapoints on the 10-point test dataset, we also evaluate these models on the same datapoint used to test the UNet3DP.
Table 9: Performance metrics for predicting $T$ with different models and datasets. Errors in [C], MSE in [C], PATs in [%] and SSIM unitless. The LGCNN-test dataset corresponds to the 1 datapoint used for testing the LGCNN approach.
Model Data Case Huber L MAE MSE PAT SSIM
UNet3DP randomK3 train 0.0020 2.4954 0.0404 0.0040 6.43 0.8281
val 0.0269 5.2901 0.1365 0.0574 38.82 0.5717
test 0.0235 4.8642 0.1314 0.0492 39.05 0.5794
UNet rK101 train 0.0010 4.2443 0.0172 0.0021 2.34 0.9960
val 0.0051 4.1972 0.0441 0.0106 12.68 0.9859
test 0.0050 4.2140 0.0426 0.0104 11.90 0.9869
* LGCNN-test 0.0048 4.3985 0.0473 0.0100 13.63 0.9827
* scaling 0.0016 4.3426 0.0202 0.0033 4.33 0.9955
2$\times$2-DDUNet rK101 train 0.0018 3.8100 0.0236 0.0038 4.47 0.9940
val 0.0079 4.0084 0.0550 0.0165 17.03 0.9825
test 0.0076 3.7006 0.0549 0.0159 17.14 0.9835
* LGCNN-test 0.0063 3.4257 0.0548 0.0128 17.42 0.9804
* scaling 0.0025 4.1806 0.0235 0.0052 6.11 0.9940
Table 10: Statistics for predicting $T$ with different models and datasets. Errors in [C], MSE in [C], PATs in [%] and SSIM unitless. The LGCNN-test dataset corresponds to the 1 datapoint used for testing the LGCNN approach. Mean $\pm$ standard deviation reported.
Model Data Case Huber L MAE MSE PAT SSIM
UNet3dp randomK3 train $0.0025\pm 0.0024$ $4.69\pm 0.17$ $0.0155\pm 0.0071$ $0.0063\pm 0.0062$ $1.92\pm 1.16$ $0.981\pm 0.012$
val $0.0219\pm 0.0008$ $4.89\pm 0.07$ $0.1114\pm 0.0023$ $0.0474\pm 0.0019$ $34.76\pm 0.99$ $0.696\pm 0.006$
test $0.0176\pm 0.0006$ $4.85\pm 0.09$ $0.1043\pm 0.0018$ $0.0368\pm 0.0015$ $34.76\pm 0.76$ $0.703\pm 0.006$
UNet101dp rK101 train $0.0011\pm 0.0002$ $4.37\pm 0.20$ $0.0182\pm 0.0020$ $0.0023\pm 0.0004$ $2.45\pm 0.27$ $0.995\pm 0.002$
val $0.0055\pm 0.0003$ $4.36\pm 0.16$ $0.0454\pm 0.0019$ $0.0114\pm 0.0006$ $12.96\pm 0.55$ $0.984\pm 0.002$
test $0.0052\pm 0.0003$ $4.35\pm 0.23$ $0.0441\pm 0.0019$ $0.0110\pm 0.0006$ $12.49\pm 0.48$ $0.985\pm 0.002$
* LGCNN-test $0.0049\pm 0.0002$ $4.30\pm 0.16$ $0.0470\pm 0.0010$ $0.0102\pm 0.0004$ $13.51\pm 0.59$ $0.983\pm 0.002$
* scaling $0.0017\pm 0.0001$ $4.48\pm 0.15$ $0.0208\pm 0.0014$ $0.0035\pm 0.0002$ $4.38\pm 0.17$ $0.995\pm 0.001$
2$\times$2-DDUNet101dp rK101 train $0.0014\pm 0.0003$ $4.11\pm 0.25$ $0.0203\pm 0.0026$ $0.00300\pm 0.0007$ $3.24\pm 0.68$ $0.995\pm 0.001$
val $0.0079\pm 0.0002$ $4.20\pm 0.25$ $0.0564\pm 0.0008$ $0.01648\pm 0.0005$ $17.32\pm 0.22$ $0.981\pm 0.002$
test $0.0075\pm 0.0001$ $4.05\pm 0.22$ $0.0552\pm 0.0008$ $0.01580\pm 0.0002$ $16.94\pm 0.36$ $0.982\pm 0.001$
* LGCNN-test $0.0057\pm 0.0003$ $4.00\pm 0.20$ $0.0526\pm 0.0015$ $0.01171\pm 0.0006$ $16.44\pm 0.63$ $0.981\pm 0.002$
* scaling $0.0025\pm 0.0001$ $4.04\pm 0.20$ $0.0251\pm 0.0007$ $0.00514\pm 0.0002$ $6.39\pm 0.17$ $0.994\pm 0.001$
LGCNN randomK3 Step 3 train $0.0001\pm 0.0001$ $2.64\pm 0.18$ $0.0064\pm 0.0008$ $0.0003\pm 0.0001$ $0.29\pm 0.08$ $0.996\pm 0.001$
val $0.0032\pm 0.0001$ $2.49\pm 0.29$ $0.0413\pm 0.0008$ $0.0065\pm 0.0003$ $10.88\pm 0.35$ $0.912\pm 0.003$
test $0.0025\pm 0.0000$ $2.74\pm 0.33$ $0.0382\pm 0.0006$ $0.0049\pm 0.0001$ $9.44\pm 0.35$ $0.918\pm 0.003$
scaling $0.0008\pm 0.0000$ $3.03\pm 0.31$ $0.0179\pm 0.0006$ $0.0016\pm 0.0001$ $2.54\pm 0.14$ $0.946\pm 0.005$
Table 11: Statistics for predicting $\vec{v}$ with the randomK dataset. Errors in [m/y], MSE in [m2/y2], SSIM unitless. Mean $\pm$ standard deviation reported.
Model Data Output Case Huber L MAE MSE SSIM
LGCNN randomK3 $v_{x}$ train 14.11 $\pm$ 11.52 96.63 $\pm$ 23.29 14.60 $\pm$ 11.53 380.19 $\pm$ 529.31 0.997 $\pm$ 0.002
$v_{y}$ train 14.30 $\pm$ 7.10 135.26 $\pm$ 20.36 14.80 $\pm$ 7.11 361.61 $\pm$ 296.21 0.991 $\pm$ 0.006
$v_{x}$ val 28.23 $\pm$ 7.76 445.13 $\pm$ 35.50 28.72 $\pm$ 7.77 1601.10 $\pm$ 587.80 0.989 $\pm$ 0.003
$v_{y}$ val 27.33 $\pm$ 2.25 343.35 $\pm$ 26.80 27.82 $\pm$ 2.25 1587.57 $\pm$ 252.10 0.982 $\pm$ 0.002
$v_{x}$ test 26.94 $\pm$ 5.99 216.74 $\pm$ 18.92 27.44 $\pm$ 5.99 1368.52 $\pm$ 431.48 0.990 $\pm$ 0.002
$v_{y}$ test 30.38 $\pm$ 6.18 249.11 $\pm$ 11.37 30.88 $\pm$ 6.18 1848.47 $\pm$ 637.10 0.982 $\pm$ 0.005
$v_{x}$ scaling 27.89 $\pm$ 6.53 286.08 $\pm$ 22.96 28.38 $\pm$ 6.53 1486.23 $\pm$ 451.52 0.990 $\pm$ 0.002
$v_{y}$ scaling 27.60 $\pm$ 2.43 328.80 $\pm$ 24.95 28.09 $\pm$ 2.43 1578.87 $\pm$ 258.66 0.979 $\pm$ 0.003
LGCNN realK $v_{x}$ train 29.66 $\pm$ 2.29 592.17 $\pm$ 37.66 30.16 $\pm$ 2.29 2326.33 $\pm$ 327.11 0.988 $\pm$ 0.005
$v_{y}$ train 28.93 $\pm$ 5.47 397.09 $\pm$ 72.78 29.43 $\pm$ 5.47 1778.28 $\pm$ 551.97 0.997 $\pm$ 0.009
$v_{x}$ val 22.17 $\pm$ 2.24 170.47 $\pm$ 9.75 22.66 $\pm$ 2.24 909.39 $\pm$ 139.94 0.988 $\pm$ 0.002
$v_{y}$ val 20.20 $\pm$ 4.07 113.33 $\pm$ 12.37 20.70 $\pm$ 4.08 645.11 $\pm$ 221.50 0.997 $\pm$ 0.001
$v_{x}$ scaling 79.33 $\pm$ 21.80 532.35 $\pm$ 140.50 79.83 $\pm$ 21.80 9404.62 $\pm$ 4048.09 0.953 $\pm$ 0.014
$v_{y}$ scaling 53.39 $\pm$ 12.64 602.41 $\pm$ 128.79 53.89 $\pm$ 12.64 4621.25 $\pm$ 1755.50 0.980 $\pm$ 0.006
LGCNN: Metrics of training and ablation study
The results of the LGCNN and DDU-Net, evaluated on both synthetic and realistic permeability fields, for the training, validation, testing, and scaling datasets are shown in Table 12 (for Step 1 - predict $\vec{v}_{\text{pred}}$) and Table 13 (for Step 3 - predict $T$). The column "data" refers to the dataset that a model was trained on and applied to, i.e., "randomK" stands for the baseline dataset 3dp, "randomK101" for 101dp. An asterisk indicates that the model was evaluated on a dataset different from the one it was trained on. Mostly relevant for the vanilla approaches that are trained on 101dp and applied to scaling and test of 3dp.
Table 12: Performance metrics for for predicting $\vec{v}$ with different models and datasets. Errors in [m/y], MSE in [m2/y2], SSIM unitless.
Model Data Output Case Huber L MAE MSE SSIM
1st Step
LGCNN randomK $v_{x}$ train 9.4732 132.6672 9.9620 171.0032 0.9972
$v_{y}$ train 11.4005 223.4601 11.8902 275.6190 0.9937
$v_{x}$ val 22.2241 343.9907 22.7179 1102.3721 0.9905
$v_{y}$ val 26.6078 274.8036 27.1026 1524.5099 0.9841
$v_{x}$ test 21.8237 190.8046 22.3178 972.5668 0.9911
$v_{y}$ test 32.2488 256.2519 32.7444 2031.3357 0.9812
* $v_{x}$ scaling 24.4314 294.0457 24.9261 1204.1154 0.9911
* $v_{y}$ scaling 25.7847 367.6891 26.2795 1463.8218 0.9820
Experiment: trained on full image
LGCNN randomK $v_{x}$ train 3.1983 39.9380 3.6714 21.5058 0.9993
$v_{y}$ train 3.0496 43.9018 3.5177 21.4905 0.9988
$v_{x}$ val 25.2418 417.1447 25.7367 1421.6547 0.9896
$v_{y}$ val 29.6294 276.8029 30.1247 1827.3134 0.9824
$v_{x}$ test 23.3646 202.6397 23.8584 1144.4479 0.9913
$v_{y}$ test 35.4289 332.9557 35.9249 2490.4104 0.9789
Replace with UNet-101dp, i.e., $1\times 1$ subdomain-DDUNet, trained on 101 datapoints
DDUNet randomK101 $v_{x}$ train 3.5468 678.7761 4.0134 86.6731 0.9767
$v_{y}$ train 4.1030 690.1506 4.5816 100.2767 0.9934
$v_{x}$ val 3.8368 647.1833 4.3058 90.3292 0.9739
$v_{y}$ val 4.3649 685.6656 4.8433 102.8601 0.9931
$v_{x}$ test 3.6643 584.8119 4.1325 83.4062 0.9731
$v_{y}$ test 4.1355 547.8865 4.6131 87.7819 0.9924
* $v_{x}$ LGCNN-test 12.2161 1047.8464 12.7036 896.2758 0.9016
* $v_{y}$ LGCNN-test 13.9033 2594.9504 14.3878 2317.6863 0.9903
* $v_{x}$ scaling 4.7412 659.8439 5.2152 79.2508 0.9699
* $v_{y}$ scaling 5.2686 456.1904 5.7492 83.9735 0.9948
Replace with 2x2-DDUNet-101dp, i.e.,2×$\times 2$ 2 subdomains
DDUNet randomK101 $v_{x}$ train 5.3928 422.4526 5.8815 57.6045 0.9580
$v_{y}$ train 4.0476 696.2919 4.5116 58.8457 0.9972
$v_{x}$ val 5.7663 471.6309 6.2545 76.4372 0.9552
$v_{y}$ val 4.5085 644.1844 4.9751 71.5679 0.9973
$v_{x}$ test 5.5847 319.9844 6.0735 59.2501 0.9526
$v_{y}$ test 4.2292 535.0656 4.6948 54.5020 0.9972
* $v_{x}$ LGCNN-test 13.1829 2453.5881 13.6735 1367.9061 0.9325
* $v_{y}$ LGCNN-test 14.6481 2029.0269 15.1332 2221.4226 0.9919
* $v_{x}$ scaling 7.6311 421.4186 8.1184 125.9511 0.9314
* $v_{y}$ scaling 6.9406 400.2490 7.4200 135.7929 0.9955
Domain Transfer, i.e., to more realistic data
LGCNN realK $v_{x}$ train 13.6890 122.7981 14.1819 340.6713 0.9973
$v_{y}$ train 9.0680 71.2576 9.5619 126.7222 0.9991
$v_{x}$ val 14.9187 106.8607 15.4095 380.3762 0.9939
$v_{y}$ val 10.1675 74.5570 10.6605 148.2475 0.9993
* $v_{x}$ scaling 109.5079 459.1620 110.0078 13570.0000 0.9462
* $v_{y}$ scaling 17.1118 240.9406 17.6051 616.5186 0.9965
Table 13: Performance metrics for predicting $T$ with different models and datasets. Errors in [°C], MSE in [°C2], PATs in [%], SSIM unitless.
Model Data Case Huber L MAE MSE PAT SSIM
3rd Step: trained and applied to $v_{sim}$
LGCNN randomK $v_{sim}$ train 1.87e-5 2.2536 0.0014 3.95e-5 0.03 0.9997
val 0.0027 2.8857 0.0369 0.0054 8.61 0.9283
test 0.0021 2.8990 0.0347 0.0041 7.54 0.9304
* scaling 0.0007 3.0250 0.0168 0.0014 2.05 0.9510
Full Pipeline, i.e., Step 3 trained on $v_{sim}$, but applied to $v_{pred}$
LGCNN randomK $v_{pred}$ train 0.0121 5.1006 0.0642 0.0272 18.17 0.8700
val 0.0188 4.2264 0.0967 0.0411 28.96 0.7625
test 0.0147 4.2120 0.0905 0.0307 28.92 0.7637
* scaling 0.0065 4.9366 0.0413 0.0141 10.87 0.8654
Experiments on 3rd Step
Trained in sequence, i.e., trained on $v_{pred}$
LGCNN randomK $v_{pred}$ train 0.0032 4.0750 0.0198 0.0068 3.73 0.9794
val 0.0195 3.8178 0.1003 0.0423 30.46 0.7442
test 0.0139 4.0804 0.0901 0.0289 29.44 0.7553
Trained with zero-padding
LGCNN randomK $v_{sim}$ train 0.0002 1.8465 0.0140 0.0004 0.33 0.9518
val 0.0032 2.9051 0.0465 0.0064 11.31 0.8686
test 0.0024 2.3391 0.0418 0.0048 9.19 0.8838
Trained on full image
LGCNN randomK $v_{sim}$ train 0.0002 3.5648 0.0022 0.0004 0.09 0.9992
val 0.0027 3.5088 0.0375 0.0055 8.81 0.9217
test 0.0021 3.6420 0.0351 0.0043 7.41 0.9245
Replace with UNet-101dp: Step 3 trained and applied on $v_{sim}$
DDUNet rK101 train 0.0019 3.8491 0.0283 0.0038 5.35 0.9911
val 0.0027 3.9381 0.0350 0.0055 8.89 0.9895
test 0.0025 3.7999 0.0343 0.0052 8.50 0.9903
* LGCNN-test 0.0032 4.4658 0.0416 0.0067 11.69 0.9835
* scaling 0.0010 4.3489 0.0176 0.0022 3.07 0.9954
Replace with 2x2-DDUNet-101dp: Step 3 trained and applied on $v_{sim}$
DDUNet rK101 train 0.0015 4.0624 0.0200 0.0032 3.44 0.9952
val 0.0040 4.1863 0.0410 0.0083 12.54 0.9904
test 0.0039 4.0807 0.0411 0.0081 12.67 0.9910
* LGCNN-test 0.0029 4.2772 0.0376 0.0058 10.96 0.9915
* scaling 0.0012 3.9946 0.0192 0.0025 4.73 0.9968
Domain Transfer, i.e. to more realistic data
3rd Step
LGCNN realK $v_{sim}$ train 0.0002 0.7704 0.0139 0.0005 0.43 2.9107
val 0.0005 0.8222 0.0175 0.0010 2.18 0.9672
* scaling 0.0004 0.8052 0.0189 0.0008 0.92 0.9497
Full Pipeline
LGCNN realK $v_{pred}$ train 0.0049 2.5437 0.0534 0.0100 17.58 2.4923
val 0.0137 2.3194 0.0841 0.0275 27.79 0.7510
* scaling 0.0022 2.0511 0.0394 0.0044 10.02 0.8708
LGCNN+random $k$: Performance of Step 1
The model generally obtains good results in Figure 3, even for cells that are far away from injection points.
Figure 6: Prediction of $\vec{v}$ with LGCNN. Left: Test datapoint with labels $v_{x}$ and $v_{y}$. Middle: Step 1 prediction for $v_{x}$, $v_{y}$. Right: Error of the $v_{x}$, $v_{y}$ prediction.
LGCNN+random $k$: Experiment on Inputs to Step 3
Additional tests show that excluding other inputs, either $i$ alone or both $i$ and $k$, raises prediction error by 58–121%. We also evaluated alternative time-stepping schemes for solving the IVP. Replacing the 5th-order implicit Runge–Kutta method with explicit 2nd- or 4th-order schemes accelerates computation, but increases prediction error by 16–23%—a moderate degradation compared to the complete removal of streamline inputs. Nonetheless, we retain the implicit scheme for its superior accuracy and stability. Quantitative and qualitative results for the predictions are shown in Figure 7 and Table 14.
Table 14: Experiment on 3rd step: Test metrics for predicting $T$ with different input combinations. Errors in [°C], MSE in [°C2], PAT in [%], SSIM unitless.
Inputs Huber L MAE MSE PAT SSIM
$ikv_{x}v_{y}$ 0.0070 2.2990 0.0712 0.0139 25.55 0.7662
$ikv_{x}v_{y}s$ 0.0041 1.8674 0.0545 0.0083 20.56 0.8368
$ikv_{x}v_{y}s_{o}$ 0.0057 2.6623 0.0598 0.0114 20.64 0.8423
$ikv_{x}v_{y}ss_{o}$ (not faded) 0.0072 2.3039 0.0744 0.0144 21.40 0.7837
$ikv_{x}v_{y}ss_{o}^{\mathrm{a}}$ 0.0031 1.8364 0.0442 0.0062 12.97 0.8828
$v_{x}v_{y}ss_{o}$ 0.0066 2.1301 0.0647 0.0132 20.64 0.8681
$kv_{x}v_{y}ss_{o}$ 0.0049 2.0925 0.0587 0.0097 20.32 0.8732
explicit RK, order 4 0.0038 3.2636 0.0486 0.0076 15.80 0.8871
explicit RK, order 2 0.0036 2.2034 0.0463 0.0072 13.75 0.8830
a new run to be comparable to the others in this experiment: trained
with Huber validation loss, hence the results differ slightly wrt. to Table 3.
Figure 7: 1st Column: Label, input $k$, 2nd-5th: Without $(s,s_{o})$, $(s_{o})$, $(s)$, not-faded streamlines $(s,s_{o})$, 6th: include all inputs, 7th-8th: Without $(i,k)$, $(i)$. Absolute errors capped at 1°C for better visualizations. Maximum errors are listed in Table 14.
LGCNN+realistic $k$: Performance of 3rd step and full pipeline
The qualitative performance is observable in Figure 8, where we see coherent streamlines and plume structures for both the isolated 3rd step and the full pipeline.
Figure 8: 1st Column: Label, input $k$. 2nd Column: 3rd step prediction of $T$ and error with respect to the label. 3rd Column: Predicted $T$ and error of the full pipeline.
Training and inference times
Table 15 summarizes the training and inference times, number of epochs, and dataset splits (train:val:test) for each of the three steps in our pipeline, both for the LGCNN trained on partitioned and full datasets. Although both approaches exhibit similar inference times, they show significant differences in training time, with the partitioned approach yielding better results.
Table 15: Training measurements on the randomK dataset.
Data Points Epochsa Training Timea Inference Time
(train:val:test)
1st Step (partitioned) 20,736:1:1 19 9.5 min 0.02 s
1st Step (full) 1:1:1 9,688 92.6 min 0.02 s
2nd Step 1:1:1 - - 9.82 s
3rd Step (partitioned) 82,944:1:1 14 31.5 min 0.03 s
3rd Step (full) 1:1:1 9,671 92.1 min 0.02 s
aEarly stopping: measurements until best validation loss.
In Table 16, the number of epochs and the total training time for the data-driven approaches are shown. For Step 1, both the UNet and DDUNet need many epochs and comparable training time to converge; however, for the third step and the full pipeline, the DDUNet significantly reduces both the number of epochs and the total training time required to reach convergence.
[…]
Epochsa Training Timea
1st Step
UNet101dp 738 5.497 hours
2$\times$2 DDUNet101dp 735 4.787 hours
3rd Step
UNet101dp 726 8.343 hours
2$\times$2 DDUNet101dp 303 3.508 hours
Full Pipeline
UNet101dp 267 3.680 hours
2$\times$2 DDUNet101dp 97 1.222 hours
aEarly stopping: measurements until best validation loss.
A.4 Hardware Specifications
The 2$\times$2 DDU-Net101dp and UNet101dp models, trained on the large data-driven dataset of 101 samples, were trained and evaluated on the DelftBlue cluster [DHPC2024], using NVIDIA V100 GPUs with 32 GB memory. All training was conducted using PyTorch 2.1.0 with CUDA 11.6 acceleration.
Training and evaluation of the LGCNN model were performed on a single NVIDIA A100-SXM4 GPU. Data generation was carried out on a dual-socket system equipped with AMD EPYC 9274F CPUs.
A.5 Glossary
A list of the most relevant physical properties used in our paper is provided in Table 17.
Table 17: Glossary of Abbreviations.
Abbr. Parameter
$t$ time
$X(t_{0})$ property X at initial time
$X(t_{\text{end}})$ property X at quasi steady-state
$X_{pred}$ predicted property X
$i$ positions of heat pumps
$Q_{inj}$ injected mass rate
$\Delta T_{inj}$ injected temperature difference
$k$ permeability
$p$ hydraulic pressure
$g=\nabla p$ hydraulic pressure gradient
$\vec{v}=\left(v_{x},v_{y}\right)$ flow velocity
$\vec{s}$ both streamline fields
$s$ central streamlines after all $i$
$s_{o}$ streamlines with transversal offset to $i$
$T$ temperature

### v2

@@
-library.bib
+Resolving Extreme Data Scarcity by Explicit Physics Integration: An Application to Groundwater Heat Transport
 Abstract
-Machine learning methods often struggle with real-world applications in science and engineering due to limited or low-quality training data.
-In this work, the example of groundwater flow with heat transport is considered; this corresponds to an advection-diffusion process under heterogeneous flow conditions, that is, spatially distributed material parameters and heat sources.
-Classical numerical simulations are costly and challenging due to high spatio-temporal resolution requirements and large domains.
-While often computationally more efficient, purely data-driven surrogate models face difficulties, particularly in predicting the advection process, which is highly sensitive to input variations and involves long-range spatial interactions.
-Therefore, in this work, a Local-Global Convolutional Neural Network (LGCNN) approach is introduced.
-It combines a lightweight numerical surrogate for the transport process (global) with convolutional neural networks for the groundwater velocity and heat diffusion processes (local).
-With the LGCNN, a city-wide subsurface temperature field is modeled, involving a heterogeneous groundwater flow field and one hundred groundwater heat pump injection points forming interacting heat plumes over long distances.
-The model is first systematically analyzed based on random subsurface input fields.
-Then, the model is trained on a handful of cut-outs from a real-world subsurface map of the Munich region in Germany, and it scales to larger cut-outs without retraining.
-All datasets, our code, and trained models are published for reproducibility.
+Real-world flow applications in complex scientific and engineering domains, such as geosciences, challenge classical simulation methods due to large spatial domains, high spatio-temporal resolution requirements, and potentially strong material heterogeneities that lead to ill-conditioning and long runtimes.
+While machine learning–based surrogate models can reduce computational cost, they typically rely on large training datasets that are often unavailable in practice.
+To address data-scarce settings, we revisit the structure of advection–diffusion problems and decompose them into multiscale processes of locally and globally dominated components, separating spatially localized interactions and long-range effects.
+We propose a Local–Global Convolutional Neural Network (LGCNN) that combines a lightweight numerical model for global transport with two convolutional neural networks addressing processes of a more local nature.
+We demonstrate the performance of our method on city-scale geothermal heat pump interaction modeling and show that, even when trained on fewer than five simulations, LGCNN generalizes to arbitrarily larger domains, and can be successfully transferred to real subsurface parameter maps from the Munich region, Germany.
+Machine Learning, ICML
 1 Introduction
-Many real-world systems, across environmental engineering, geoscience, and biomedicine, feature transport phenomena governed by coupled advection–diffusion processes, e.g., pollutant transport through air or water, chemical reactions, or heat transport in flowing media.
-In such settings, interactions can span large distances, long timescales and many components.
-Datasets based on physical measurements are often limited by cost, logistics, or experimental constraints.
-While numerical simulations can address these limitations to some extent, they become impractically slow or computationally expensive as domain size or resolution increases.
-Machine learning (ML) models offer fast predictions but typically require large training datasets and often struggle to scale to larger domains.
-This trade-off underscores the need for data-efficient surrogates that can capture both local and long-range interactions, and scale to arbitrary domain sizes and numbers of interacting components.
-To this end, we introduce the Local–Global Convolutional Neural Network (LGCNN), a scalable, physics-inspired approach that leverages the structure of advection–diffusion problems by decoupling local and global transport processes.
-We apply this approach to the climate change-relevant application of modeling city-wide subsurface temperature fields driven by heat and mass injections of dozens of open-loop groundwater heat pumps (GWHP)[pophillat2020impact, heatpumps, hahnlein2013] in the region of Munich, Germany[geokw22].
-A city-wide optimization of the positions of all GWHPs requires calling a prediction model dozens of times with slightly adapted GWHP positions to minimize negative interactions between different heat plumes and pumps.
-For this scenario, our model should predict the temperature field $T$ in an arbitrarily large domain based on the inputs of the subsurface parameters of a hydraulic pressure gradient $\nabla p$ and a heterogeneous subsurface permeability field $k$, and arbitrary heat pump locations $i$.
-Labels are generated by expensive simulations, but only for a small number of cases and entirely offline.
-As sketched in Figure 1, the LGCNN uses CNNs where they excel, and simple numerical surrogates where CNNs struggle: a first CNN predicts the velocities $\vec{v}$ from local relations of the inputs $p$, $k$ and $i$ (Step 1); a fast numerical component calculates the streamlines $\vec{s}$ based on this (Step 2); a second CNN calculates the temperature distribution $T$ in the whole domain based on all prior inputs and outputs (Step 3).
-Since Step 2 covers all non-local effects, Step 3 can again take advantage of CNNs to learn local patterns.
+Many learning problems in scientific and engineering domains (Sharma et al., 2021; Sarker, 2021; Angra & Ahuja, 2017; Jhaveri et al., 2022) are governed by underlying physical processes.
+Yet they challenge classical numerical simulations due to large spatial domains, high resolution requirements, and potential ill-conditioning.
+In addition, many applications require repeated evaluations, for example for uncertainty quantification, design optimization, and real-time control.
+From a machine learning (ML) perspective, such problems are attractive targets for surrogate modeling, as trained models can dramatically reduce inference cost, even if offline training cost is substantial.
+However, in real-world advection-diffusion problems, standard ML approaches face two central difficulties (Dietrich & Schilders, 2025).
+First, training data are often scarce: measurements are indirect or sparse, and generating labeled data via high-fidelity simulations is computationally expensive.
+Second, the underlying dynamics exhibit multiscale behavior with both short-range local interactions and long-range global dependencies, making it difficult for data-driven models to generalize in data-scarce regimes.
+Convolutional architectures such as UNets (Ronneberger et al., 2015a) effectively model local spatial patterns (Thuerey et al., 2020; Jhaveri et al., 2022), but representing non-local interactions requires deep receptive fields and large training datasets.
+Operator learning approaches, such as Fourier Neural Operators (FNOs) (Li et al., 2020), address long-range dependencies more explicitly in the frequency domain (Choi et al., 2024), but inherently introduce additional assumptions, including periodic boundary conditions, homogeneous coefficients, and fixed domain sizes, which limit their applicability in many real-world settings.
+Physics-informed ML aims to mitigate data scarcity by embedding physical constraints or priors into the learning process.
+Physics-Informed NNs (PINNs) (Dissanayake & Phan-Thien, 1994; Lagaris et al., 1998; Raissi et al., 2019), for instance, replace or augment data losses with squared differential equation residuals, enabling learning with limited labeled data.
+Despite notable progress (Cuomo et al., 2022; Rao et al., 2020; Sun et al., 2020; Grimm et al., 2025; Wen et al., 2022; Cai et al., 2021), such methods remain at an early stage of methodological maturity and often struggle in complex, high-dimensional scenarios, for example by converging to trivial solutions (Krishnapriyan et al., 2021).
+In our approach, rather than increasing architectural complexity, we build on a simple, well-suited base architecture and exploit the physical structure, i.e., of the underlying advection–diffusion system, explicitly.
+Such systems naturally decompose into processes governed by predominantly local interactions, such as diffusion acting at neighborhood scales and inferable from local information, as well as global transport effects, such as long-range advection, that require propagating information across large parts of the domain.
+Building on this observation, we propose the Local–Global CNN (LGCNN), a modular architecture that separates local learning from explicit global transport modeling.
+LGCNN combines CNNs that are well-suited for locally dominant, diffusion-driven processes with a lightweight numerical surrogate that captures long-range transport effects.
+By inserting this inexpensive, physically motivated component at a critical point in the inference pipeline, LGCNN captures long-range dependencies without relying on deep architectures, large receptive fields, or extensive training data.
+In contrast to physics-informed approaches that softly impose physical constraints through loss terms, such as PINNs, LGCNN handles global transport at inference time via a dedicated, non-learnable module, leading to improved data efficiency and robustness.
+We evaluate LGCNN on the task of predicting city-scale subsurface temperature fields induced by one hundred interacting groundwater heat pumps in the Munich region, Germany (Zosseder et al., 2022).
+This problem provides a challenging testbed for data-efficient learning, as the temperature field depends on heterogeneous subsurface parameters, arbitrary source configurations, and long-range advective transport over several kilometers and decades.
+We show that, even when trained on fewer than five simulated samples, LGCNN generalizes to much larger spatial domains without retraining.
+Our approach also transfers from synthetic inputs to real subsurface parameter maps derived from borehole measurements.
 Figure 1: Left: Map of $k$ of whole region of Munich.
 Right: Schematics of our Local-Global CNN-based approach (LGCNN) with 3 physics-inspired steps: CNN ($pki\rightarrow\vec{v}$), simplified solver ($i\vec{v}\rightarrow\vec{s}$), CNN ($pki\vec{v}\vec{s}\rightarrow T$).
-We outline related work in Section 2, then define our datasets and metrics in Section 3.
-In Section 4, we benchmark purely data-driven models and show their performance and limitations on this low-data challenge.
-To address data scarcity, we introduce LGCNN in Section 5.
-Section 6 emphasizes generalization and scalability from a single datapoint, while Section 7 demonstrates transfer to real-world subsurface parameters extracted from measurement maps such as in Figure 1.
+Formally, the learning task is to predict a temperature field $T$ over arbitrarily large domains from heterogeneous permeability fields $k$, hydraulic pressure gradients $\nabla p$, and heat pump locations $i$.
+Training labels are generated offline using expensive numerical simulations for only a small number of configurations.
+As illustrated in Figure˜1, LGCNN addresses this setting via a three-stage pipeline: (1) a CNN predicts the velocity field $\vec{v}$ from local relations between $p$, $k$, and $i$; (2) a cheap numerical surrogate computes streamlines $\vec{s}$ to model global transport; and (3) a second CNN predicts the temperature field $T$ conditioned on inputs and intermediate representations.
+This explicit separation of local learning and global transport is key to the scalability of the approach.
+In groundwater heat pump (GWHP) modeling, existing work (Davis et al., 2023; Pelzer & Schulte, 2024; Scheurer, 2021) focuses on isolated or pairwise pump interactions, often using UNet-based architectures with or without physics-based regularization.
+These methods typically require substantial training data and have not been demonstrated to scale to city-wide domains with many interacting sources and heterogeneous aquifers.
 Contributions
-Our method strongly reduces data requirements for modeling complex scenarios, increases reliability, and ensures scalability to larger domains by a domain-specific modular design of physics-aided machine learning for an input-sensitive scenario with long-distance effects.
-Subsurface temperature prediction serves as our example application, with the approach expected to generalize to related domains.
+We introduce a modular, physics-inspired architecture that captures complex, multiscale dynamics by combining local convolutional feature learning with explicit long-range transport modeling.
+To avoid deep architectures and large datasets for modeling nonlocal interactions, LGCNN enforces global coupling through inexpensive numerical transport modules.
+This structural inductive bias enables inherent generalization and scalable inference across domains and problem sizes from very limited data, i.e., less than five simulations.
+Compared to classical high-fidelity simulations, our approach achieves an inference-time speedup of approximately 2,000 times.
 Limitations
-The current approach is limited to two-dimensional and steady-state datasets.
-2 Related Work
-Deep learning surrogates in scientific modeling can be broadly categorized into purely data-driven architectures, physics-informed or domain-structured models, and highly problem-specific hybrid approaches.
-In this section, we summarize the developments in each category and existing efforts in GWHP modeling.
-Convolutional Neural Networks (CNNs), particularly UNet [CNN15] variants, are widely used for spatial prediction tasks due to their locality-aware architecture of moving-window-like kernels.
-Applications include flow around airfoils [thuerey_airfoils] and various engineering systems [Jhaveri2022A].
-For general reviews on ML in scientific and real-world engineering contexts, see [sharma2021machine, sarker2021machine, angra2017machine].
-To improve generalization and reduce data requirements, domain-structured models embed physical priors or symmetries into their architecture.
-This includes Fourier Neural Operators (FNOs) for efficient multi-scale modeling [fno, choi2024applications], thermodynamics-preserving networks [hernandez2021structure], and rotation-equivariant CNNs (ECNNs) [weiler2023ECNNs].
-Physics-Informed Neural Networks (PINNs) go further by replacing data loss with physics-based constraints, allowing for training without labeled data [PINN19, cuomo2022scientific], and have been applied to fluid dynamics and inverse problems [pinn_rao, pinn_sun, cai_physics-informed_2021].
-Hybrid approaches combine architectures and integrate domain or physics knowledge to improve general models, for instance, in low-data regimes.
-Examples include staged CNNs for atmospheric plume dispersion [fernandez2024staged], physics-informed FNOs for traffic flow [thodi2024fourier], CNN–MLP hybrids for optimizing positions of oil wells [yousefzadeh2025determination], and physics-guided CNNs for seismic response prediction [zhang2020physics].
-Deep Learning for GWHP Modeling
-In the context of groundwater modeling with heat transport, recent work focuses on isolated or pairwise interacting GWHPs.
-Most use UNet-based architectures, optionally with physics-loss terms [davis2023deep, pelzer2024, Stefania].
-Although effective in simple, homogeneous aquifers, these models rely on large training datasets or simulated inputs and have not been scaled to city-wide domains with many pumps.
-UNets are the dominant architecture because of their strong performance on spatial data, making them the baseline for comparison.
-We briefly evaluated PINNs and FNOs as alternatives.
-PINNs, while promising for low-data fluid dynamics tasks [pinn_rao, pinn_sun, cai_physics-informed_2021, pdebench], struggled with complex scenarios (cf.
-[krishnapriyan2021characterizing]) such as heterogeneous media and discontinuous source terms in preliminary tests.
-FNOs, though capable of modeling local and global dependencies, struggled with large domain sizes due to high memory requirements.
-Additionally, they performed poorly in the presence of multiple sources, consistent with [liu2024neural].
+The current implementation targets a two-dimensional steady-state setting and assumes weak or spatially localized temperature-induced flow perturbations.
+While this work is restricted to this regime, the proposed framework is easily extendable to three-dimensional and transient problems.
+Empirical validation of these extensions is left to future work due to the lack of suitable training data.
 2 Datasets and Metrics
 Datasets
-Inputs for the neural networks consist of a heterogeneous permeability field $k$, an initial hydraulic pressure field $p$, and a one-hot-encoded field of heat pump positions $i$.
-The (interim) labels of velocity $\vec{v}$ and temperature $T$ fields are simulated, in our case with Pflotran [pflotran], until a quasi-steady state is reached after $\approx$27.5 years simulated time [bawue_arbeitshilfe].
-All data are normalized to $[0,1]$, and stored in PyTorch format for training.
-We generate two types of datasets, one is based on synthetic, the other on real permeability fields $k$.
-Both cover a $12.8\times 12.8$ km2 domain with 2 560×2 $\times 2\,560$ 2 560 cells.
-The baseline dataset uses Perlin noise [perlin] to generate random, heterogeneous $k$.
-Three simulations (3dp) with different fields for $i$ and $k$ are run, generating one datapoint each for training, validation, and testing, plus a simulation 4$\times$ larger to assess scalability.
-For training vanilla neural networks in Section 4, we generate an additional dataset of 101 datapoints (101dp), split into 73:18:10 for train:val:test.
-Runtimes are $\approx$27 hours per simulation and 123 hours for the larger domain.
-The more realistic dataset builds on $k$ fields cut from maps of borehole measurements in the Munich region[lfu_gepo].
-This dataset consists of four simulations (three for training and one for validation) and one larger simulation for testing scalability.
-Due to constraints of the available measurements, see Figure 1, the large-scale simulation extends only in length, resulting in a rectangular domain twice the length.
+Inputs for the NNs consist of a heterogeneous permeability field $k$, an initial hydraulic pressure field $p$, and a one-hot-encoded field of heat pump positions $i$ of a constant injection rate.
+The (interim) labels of velocity $\vec{v}$ and temperature $T$ fields are simulated with Pflotran (Lichtner et al., 2015a) until a quasi-steady state is reached after $\approx$27.5 years simulated time (Umweltministerium Baden-Württemberg, 2009).
+[…]
+The more realistic dataset builds on $k$ fields cut from maps of borehole measurements in the Munich region, Germany (Bayerisches Landesamt für Umwelt, 2015).
+This dataset consists of four simulations (4dp) (three for training and one for validation) and one larger simulation for testing scalability.
+Due to constraints of availability of subsurface measurements, see Figure˜1, the large-scale simulation is on a rectangular domain of only twice the length of the training data.
 Runtimes range from 38 to 91 hours (average 58 hours), and 134 hours for the scaling-test domain.
-Variation across the dataset stems from different heat pump placements in $i$ and from the specific regions where $k$ is extract from the available measurement domain.
-For detailed information on simulation procedures and hydro-geological parameters, see Section A.1; for hardware specifications see Section A.4.
+Variation within the dataset for training stems from different heat pump placements in $i$ and from the location where $k$ is extract from the Munich region in 1.
+Information on simulations and hydro-geological parameters are in Section˜A.1; hardware specifications in Section˜A.4.
 Performance Metrics
-We evaluate model accuracy per output dimension ($v_{x},v_{y},T$) separately using Mean Absolute Error (MAE) [naser2023error], Mean Squared Error (MSE) [naser2023error], Maximum Absolute Error ($L_{\infty}$) Structural Similarity Index Measure (SSIM) [ssim], and application-driven metrics of Percentage Above Threshold (PAT) and Visual Assessment.
-PAT measures the percentage of cells where the absolute error of the predicted temperature exceeds the threshold of $0.1$C, corresponding to measurement precision [kabellichtlot], and is only applicable to temperature predictions.
+We evaluate model accuracy separately for each output field (directional velocities $v_{x},v_{y}$, and temperature $T$) using standard ML metrics of Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), and Huber loss; the structure-focused metric of Structural Similarity Index Measure (SSIM) (Wang et al., 2004), and application-driven metrics of Measure of Connectivity (MoC) and visual assessment.
+MoC expresses how physically reasonable the temperature fields are by measuring the percentage of predicted heat plume cells that are not connected via 4-connectivity [^1] This means axis-aligned neighbors only, excluding diagonal connections.
+to a heat pump based on a temperature threshold of $0.1$ °C, corresponding to measurement precision (UKB System Technology, 2024).
 All metrics other than SSIM are applied after re-normalization to the original data ranges to obtain physically meaningful results in [°C] or [meters/year].
-4 Application of Purely Data-Driven Approaches
-In this section, we present the performance of two purely data-driven approaches for predicting $T$ from inputs of $p,k,i$: a UNet [CNN15] and the domain-decomposition-based UNet variant, called DDUNet [DDUNet].
-Methods
-The fully convolutional vanilla UNet has an encoder-decoder architecture with skip connections between layers of equal spatial dimensions.
-Convolutional networks excel in extracting local dependencies from spatial data with locally repeating patterns.
-Hence, in theory, one datapoint covering a large domain could be enough to generalize, if only local dependencies occur.
-Global patterns would require a large receptive field and, hence, more data.
-The DDUNet is designed to handle large-scale domains under GPU memory constraints.
-It combines a UNet architecture with domain decomposition principles by partitioning the spatial domain into subdomains that are distributed across multiple GPUs.
-We use a decomposition into $2\times 2$ subdomains.
-[^1] Conceptually, DDUNet reduces to a standard UNet when trained with a single subdomain.
-Global context is communicated between subdomains via coarse feature maps through a learned communication network, maintaining spatial consistency and providing a large receptive field while reducing memory demands.
-The hyperparameters of the models are optimized according to Section A.2.1.
-Experiments
-Experiments for the vanilla UNet are conducted on the baseline datasets with 3dp and 101dp, those with DDUNet directly on the 101dp dataset, see Section A.3.
-Both models are applied to the test of 3dp and to the scaling datapoint, for direct comparability to LGCNN.
-UNet exhibits poor performance when trained on the limited 3dp dataset (UNet3dp), as demonstrated in both Figure 4 and Table 1.
-Similar performance would be expected for the DDUNet when applied to this sparse dataset.
-When trained on 101dp, however, both demonstrate robust predictive and scaling capabilities.
-The advantage of DDUNet over vanilla UNet is a reduction of memory usage and runtime, see section A.3.
+3 Application of Established ML Approaches
+To motivate a structure-specific architecture, we present the performance of six established ML approaches for predicting $T$ from $p,k,i$: a UNet, a domain decomposition-based UNet (DDUNet) (Verburg et al., 2025), a dilated UNet, UNet++ (Zhou et al., 2018), an FNO and a PINN (results in Tables˜1 and 2, parameters in Section˜A.2.3).
+PINNs and FNOs
+PINNs, while promising for low-data fluid dynamics tasks (Rao et al., 2020; Sun et al., 2020; Cai et al., 2021; Takamoto et al., 2022), struggle in preliminary tests with complex scenarios (Krishnapriyan et al., 2021), such as heterogeneous media and discontinuous source terms.
+Physics-informed CNNs (PI-CNNs) have been evaluated in simplified single-plume aquifer settings (Hirche, 2023; Pelzer & Piller, 2024) and consistently failed to outperform standard CNNs, even when the physics-based loss was used only as an auxiliary regularization term alongside a data loss.
+The largest errors occur near injection locations, likely due to discontinuous source term fields.
+The largest errors occur near injection locations, likely due to discontinuous source term fields.
+The largest errors occur near injection locations, likely due to discontinuous source term fields.
+In our 3dp problem, overfitting a PI-CNN to a single datapoint using only energy-based loss terms (cf.
+Section˜A.2.2) derived from the governing equations of the simulations fails to model $T$ but instead predicts a field that resembles $v_{y}$ (Figure˜2).
+This behavior is consistent with the model capturing only a preliminary stage of the physical process and indicates a representational limitation of the learning setup regarding transport modeling.
+While FNOs are in principle capable of modeling both short- and long-range interactions, we found them impractical for the target problem due to infeasible memory requirements at realistic domain sizes and their sensitivity to non-periodic boundary conditions.
+Additionally, they performed poorly in the presence of multiple sources, consistent with Liu-Schiaffini et al.
+(2024).
+On our example, we had to reduce the domain size to $\times480$ cells to fit it on a graphics processing unit (GPU) (Section˜A.4) during training and still only managed to predict low frequency aspects, i.e., the background temperature, in our overfitting test to one datapoint.
+UNet-based Methods
+Convolutional architectures such as UNet primarily capture local spatial dependencies and can, in principle, generalize from a single large-domain sample when interactions are local.
+Modeling global effects, however, requires enlarged receptive fields or architectural adaptations, such as dilated convolutions.
+UNet++ fuses multiscale features through nested skip connections, while still relying on convolutional locality.
+DDUNet addresses large domains under GPU memory constraints by decomposing the domain into, here $2\times 2$, subdomains with learned coarse-scale feature exchange, enabling effective global context at reduced memory cost.
+The hyperparameters of all models are optimized according to Section˜A.2.3.
+Table 1: Metrics of all experiments.
+Results for standard models and LGCNN (end-to-end and Steps 1 and 3) on synthetic and real permeability fields $k$, including ablation variants.
+Models are trained on 3dp, 101dp, or 4dp.
+Metrics are reported in physical units ([°C] for Step 3 and end-to-end, [m/y] for Step 1), except MoC ([%]) and SSIM (unitless).
+Velocity experiments are shaded due to differing scales; experiment details are referenced in column Sec.
+Only contains representative runs, statistical errors are reported in the appendix.
+Architecture Training Data Applied to Step MAE RMSE Huber SSIM MoC Sec.
+Standard UNet synth. 101dp test end-to-end 0.0473 0.1001 0.0048 0.8292 0.09 3
+UNet++ synth. 101dp test end-to-end 0.0918 0.1435 0.0102 0.5331 4.06 3
+DDUNet synth. 101dp test end-to-end 0.0548 0.1134 0.0063 0.8120 0.59 3
+UNet synth. 3dp test end-to-end 0.1908 0.6190 0.0736 0.6671 10.85 3
+Dilated UNet synth. 3dp test end-to-end 0.1200 0.2236 0.0237 0.5004 10.95 3
+PI-CNN (overfit) synth. 1dp train end-to-end 0.7532 0.8593 0.3454 0.1793 - 3
+Ours LGCNN synth. 3dp test end-to-end 0.0916 0.1738 0.0146 0.6841 0.09 5.1
+UNet synth. 3dp test 3 0.0417 0.0762 0.0029 0.8540 0.07 5.1
+UNet synth. 3dp test 1 (vx) 22.3178 31.1860 21.8237 0.9739 - 5.1
+UNet synth. 3dp test 1 (vy) 32.7444 45.0703 32.2488 0.9733 - 5.1
+Experiments UNet, in sequence synth. 3dp test 3 0.0919 0.1695 0.0139 0.6714 1.30 5.2
+UNet, w/ zero-padding synth. 3dp test 3 0.0487 0.0795 0.0031 0.7784 0.57 5.2
+UNet, w/o partitioning synth. 3dp test 3 0.0429 0.0793 0.0031 0.8349 0.29 5.2
+UNet synth. 101dp test 3 0.0416 0.0816 0.0032 0.8016 0.47 5.2
+DDUNet synth. 101dp test 3 0.0376 0.0764 0.0029 0.8843 0.26 5.2
+UNet, vary inputs synth. 3dp test 3 cf. A.3 cf. A.3 cf. A.3 cf. A.3 cf. A.3 A.3
+UNet, w/o partitioning synth. 3dp test 1 (vx) 23.8584 33.8297 23.3646 0.9765 - 5.2
+UNet, w/o partitioning synth. 3dp test 1 (vy) 35.9249 49.9040 35.4289 0.9721 - 5.2
+UNet synth. 101dp test 1 (vx) 12.7036 29.9379 12.2161 0.9437 - 5.2
+UNet synth. 101dp test 1 (vy) 324.7751 328.4992 324.2751 0.9572 - 5.2
+DDUNet synth. 101dp test 1 (vx) 13.6736 36.9852 13.1829 0.9468 - 5.2
+DDUNet synth. 101dp test 1 (vy) 322.0676 325.6697 321.5676 0.9741 - 5.2
+Scaling DDUNet (vanilla) synth. 101dp scaling end-to-end 0.0235 0.0723 0.0025 0.8874 1.18 3
+UNet (vanilla) synth. 101dp scaling end-to-end 0.0202 0.0578 0.0016 0.8914 0.54 3
+LGCNN synth. 3dp scaling end-to-end 0.0413 0.1189 0.0065 0.6511 0.06 5.1
+UNet synth. 3dp scaling 3 0.0168 0.0373 0.0007 0.9405 0.08 5.1
+UNet synth. 3dp scaling 1 (vx) 24.9261 34.7004 24.4314 0.9869 - 5.1
+UNet synth. 3dp scaling 1 (vy) 26.2795 38.2599 25.7847 0.9932 - 5.1
+Domain transfer LGCNN real 4dp val end-to-end 0.0841 0.1659 0.0137 0.6511 1.65 5.3
+UNet real 4dp val 3 0.0175 0.0319 0.0005 0.9405 3.30 5.3
+UNet real 4dp val 1 (vx) 13.3316 16.5833 12.8383 0.9869 - 5.3
+UNet real 4dp val 1 (vy) 9.0224 10.5612 8.5327 0.9932 - 5.3
+LGCNN real 4dp scaling end-to-end 0.0394 0.0665 0.0022 0.6755 3.50 5.3
+UNet real 4dp scaling 3 0.0189 0.0287 0.0004 0.8942 1.77 5.3
+UNet real 4dp scaling 1 (vx) 110.0078 116.4903 109.5079 0.9436 - 5.3
+UNet real 4dp scaling 1 (vy) 17.6051 24.8298 17.1118 0.9905 - 5.3
+UNet and dilated UNet are trained on 3dp; UNet, UNet++ and DDUNet are trained on 101dp directly (Section˜A.3).
+All models are evaluated on the 3dp test data for direct comparability to LGCNN, some on the scaling datapoint.
+Training on 3dp leads to short, disconnected plumes (Figure˜2) with high MoC.
+When trained on 101dp, UNet and DDUNet show robust predictions and reasonable scaling, with DDUNet occasionally hallucinating plumes but offering lower memory usage and shorter runtimes (see Table˜12).
+UNet++ fails due to memory constraints in the scaling test.
+Overall, well-chosen established approaches can perform reasonably (best is UNet101dp: MAE = 0.05°C, MoC = 0.09) and some can scale to larger domains, but none performs reasonably well in the data-scarce setting we are interested in (best is UNet3dp: MAE = 0.19°C, MoC = 10.85).
+Their limits are clear: they are biased toward local patterns, require large datasets to capture long-range transport, and quickly run into compute and memory constraints at realistic scales.
+In low-data regimes they yield short, disconnected plumes or hallucinated structures, and scalability can break down (e.g., FNO and UNet++ fail under memory constraints).
 […]
-Table 1: Performance metrics for predicting $T$ from $pki$ with UNet and DDUNet. Errors in [°C], MSE in [°C2], PAT in [%], and SSIM unitless. Inferred on the test of 3dp and the scaling datapoint.
-Model Test Scaling
-L MAE MSE PAT SSIM L MAE MSE PAT SSIM
-UNet3dp 4.8642 0.1314 0.0492 39.05 0.5794
-UNet101dp 4.3985 0.0473 0.0100 13.63 0.9827 4.3426 0.0202 0.0033 4.33 0.9955
-$2\times 2$-DDUNet101dp 3.4257 0.0548 0.0128 17.42 0.9804 4.1806 0.0235 0.0052 6.11 0.9940
 4 Methodology
-In data-scarce settings, performance can improve when the structure of a model reflects the underlying physics, cf.
-Section 2.
-This section hence outlines the relevant physical processes, motivates our LGCNN method, and discusses a few details of LGCNN’s individual components.
+When generic ML methods fail in data-scarce settings, performance can be improved by adopting a more task-specific approach whose structure reflects the underlying physics (Fernández-Godino et al., 2024; Thodi et al., 2024; Zhang et al., 2020; Yousefzadeh et al., 2025; Bertels & Willems, 2023).
+Accordingly, this section outlines the relevant physical processes that motivate the proposed LGCNN method and discusses key aspects of its components.
 Physics of Groundwater Flow with Heat Transport
-Transport of heat induced by heat injections of GWHPs in the subsurface is an advection–diffusion process.
-Since the Péclet number, i.e., the advection-diffusion ratio, is high in our application (Pemuch-$\gg$1, more in Section A.1), the system is advection-dominated.
-Advection is largely governed by the global hydraulic pressure gradient, driving flow from higher to lower regions.
-Locally, the flow paths are influenced by the spatial distribution of permeability $k$, cf.
-Figure 2.
-As a result, small changes in $k$ can lead to significant differences in flow paths further downstream, demonstrating high input sensitivity.
+[…]
 One-Way Coupled Approach: Local-Global CNN (LGCNN)
-In a fully coupled physical model, flow field computation and heat transport along streamlines [streamline] starting at heat pumps (advection) and their widening (diffusion) must be solved monolithically.
-By splitting effects into local and global and working with steady-state simulated $\vec{v}$ during training, we can reduce the coupling to one direction without significant loss in accuracy.
-Our simplified physical pipeline consists of three steps: (1) compute a steady-state flow field $\vec{v}$ from initial subsurface parameters $p$ and $k$, with $i$ encoding mass influx around GWHPs; (2) transport injected heat along streamlines governed by $\vec{v}$ until quasi-steady state[bawue_arbeitshilfe]; and (3) apply plume widening to these heat paths, informed by soil diffusivity and $\vec{v}$, to approximate diffusion effects.
-The resulting one-way coupled LGCNN approach (see Figure 1) can formally be described as:
-  1. Step 1
-    Velocities (local) We employ a CNN to predict the velocity field $\vec{v}=\left(v_{x},v_{y}\right)$ from $p$, $k$, and $i$:
-    $\text{CNN}\left(p,k,i\right)=\vec{v}.$ (1)
-  2. Step 2
-    Streamlines (global) Based on the predicted velocities $\vec{v}$ from Step 1, we compute streamlines $\vec{s}$ originating from all pump locations in $i$ with an initial value problem (IVP) solver:
-    $\text{IVP}\left(i,\vec{v}\right)=\vec{s}.$ (2)
-  3. Step 3
-    Temperature (local) A second CNN predicts the temperature field $T$ from the initial inputs $p$, $k$, $i$, the predicted velocities $\vec{v}$ of Step 1 and streamlines $\vec{s}$ calculated in Step 2:
-    $\text{CNN}\left(p,k,i,\vec{v},\vec{s}\right)=T.$ (3)
-The model outputs the steady-state temperature field $T$ and, using the same inputs $p$, $k$, and $i$ as a simulation, serves as its direct surrogate.
+In a fully coupled simulation, flow field computation and heat transport along streamlines (str, 2011) starting at heat pumps (advection) and their widening (diffusion) influence each other.
+Therefore, convergence requires that two-way coupling conditions are satisfied, whether using a monolithic or a segregated approach.
+By working with steady-state simulated $\vec{v}$ as interim outputs during training (cf.
+Section˜A.1 Reasoning for Decoupling Assumptions), we can split the processes into local (diffusion) and global (advection) dominated ones and hence extract three subsequent steps that are coupled in only one direction.
+Our simplified physical pipeline consists of three steps: (1) compute a steady-state flow field $\vec{v}$ from initial subsurface parameters $p$ and $k$, with $i$ encoding mass influx around GWHPs; (2) transport injected heat along streamlines governed by $\vec{v}$ until quasi-steady state (Umweltministerium Baden-Württemberg, 2009); and (3) apply plume widening to these heat paths, informed by soil diffusivity and $\vec{v}$, to approximate diffusion effects.
+The resulting one-way coupled application of LGCNN (see Figure˜1) can formally be described as:
+  1. Step 1:
+    Velocities (local) We employ a CNN to predict the velocity field $\vec{v}=(v_{x},v_{y})$ from $p$, $k$, and $i$:
+    $\text{CNN}(p,k,i)=\vec{v}.$ (1)
+  2. Step 2:
+    Streamlines (global) Based on the predicted velocities $\vec{v}$, we compute streamlines $\vec{s}$ originating from all pumps in $i$ with an initial value problem (IVP) solver:
+    $\text{IVP}(i,\vec{v})=\vec{s}.$ (2)
+  3. Step 3:
+    Temperature (local) A second CNN predicts the temperature field $T$:
+    $\text{CNN}(k,i,\vec{v},\vec{s})=T.$ (3)
+The model outputs a steady-state temperature field $T$ using the same inputs as a simulation, hence serving as its direct surrogate.
 CNN Models in the Local Steps
-The local Steps 1 and 3 are approximated by a UNet[CNN15, pelzer2024], see Section A.2 for details of our architecture.
-It is beneficial to omit zero-padding to enforce shift invariance and reliance on purely local information; this approach is also validated by Islam et al.
-in [Islam2020How].
-Due to the input sensitivity of the entire problem, each step is trained independently on simulated input data.
-In particular, the CNN in Step 3 is trained on the simulated velocities $\vec{v}_{\text{sim}}$ and streamlines based on these fields.
-Only during inference, all steps are applied sequentially, i.e., Step 3 uses the outputs of Steps 1 and 2, including the predicted $\vec{v}_{\text{pred}}$ as inputs.
+The local Steps 1 and 3 are approximated by a UNet, same as in other groundwater applications (Davis et al., 2023; Pelzer & Schulte, 2024), see Section˜A.2.1 for details of our architecture.
+It is beneficial to omit zero-padding to enforce shift invariance and reliance on purely local information (Islam et al., 2020).
+Due to the input sensitivity of the entire problem, the CNN in Step 3 is trained on simulated velocities $\vec{v}_{\text{sim}}$ and streamlines based on these fields.
+Only during inference, all steps are applied sequentially, i.e., Step 3 uses the outputs of Steps 1 and 2, e.g., by including the predicted $\vec{v}_{\text{pred}}$ as inputs.
 The training is accelerated by partitioning the data into overlapping patches, which increases the effective number of datapoints while reducing their spatial size.
-Combined with the streamlines’ embedding and localizing of global flow patterns for the fully convolutional (locally) acting CNN, this training data enrichment allows the model to train effectively on very little data, sometimes as little as a single simulation run, while generalizing to unseen and even to larger domains.
+Combined with the streamlines’ embedding and localizing of global flow patterns for the fully convolutional (locally) acting CNN, this training data enrichment allows the model to train effectively on very little data, partially as little as a single simulation run, while generalizing to unseen and even to larger domains.
 Validation and testing are conducted on independent simulation runs, with the full domain being processed as a whole instead of patches during inference.
-The architecture, input choices, and training hyperparameters — including the patch size and overlap in the datapoint partitioning — were optimized through a combination of manual and automated tree-based search using Optuna [optuna_2019]; see Section A.2.1 for details.
+The architecture, input choices, and training hyperparameters, including patch size and overlap in the datapoint partitioning, were optimized; see Section˜A.2.3 for details.
 Streamline Calculation and 2D-Embedding in the Global Step
@@
 with a lightweight numerical solver for each heat pump, where $y_{0}$ represents the location of a heat pump in $i$.
-We employ the solve_ivp function from SciPy [scipy] with an implicit fifth-order Runge-Kutta method [scipyRadaux2014] and 10 0001000010\,00010 000 time steps of one day, corresponding to the simulation time of forming a heat plume.
-The computed streamlines $y$ represent a sequence of positions in 2D, that a particle - injected at a heat pump’s location - would traverse based on $\vec{v}$.
-The corresponding grid cells are assigned values that fade linearly from 1 to 0, reflecting the time it takes to reach each position, similar to a soft one-hot encoding.
-To address the sensitivity of this process with respect to $k$ and $\vec{v}$, we calculate a set of outer streamlines $s_{o}$.
-For this, we perturb each heat pump’s location $y_{0}$ by 10 cells orthogonal to the global flow direction, i.e., $\nabla p$, and compute the corresponding streamlines.
-By incorporating $\vec{s}=\left(s,s_{o}\right)$, the subsequent Step 3 obtains a spatial, localized representation of flow paths.
+We employ an implicit fifth-order Runge–Kutta method (sci, ) with 10,000 one-day time steps, corresponding to the physical timescale of heat plume formation.
+The computed streamlines $y$ represent sequences of 2D positions traversed by a particle injected at $y_{0}$ under the velocity field $\vec{v}$.
+These paths are embedded into 2D fields along the traversed grid cells by assigning values that decay linearly from 1 to 0, reflecting the time required to reach each cell and yielding a time-weighted 2D-embedding.
+To address the sensitivity of this process with respect to $k$ and $\vec{v}$, we calculate the outer streamlines $s_{o}$.
+For this, we perturb each heat pump’s location $y_{0}$ by $\pm$10 cells orthogonal to the global flow direction, i.e., $\nabla p$, and compute the corresponding streamlines.
+By incorporating $\vec{s}=\left(s,s_{o}\right)$ as two inputs to Step 3, it obtains a spatial, localized representation of flow paths.
 In future work, we aim to explore probabilistic perturbations to compute the mean and standard deviation of streamlines, which, while being computationally more expensive, could be efficiently parallelized on GPUs.
-6 Results on Synthetic Permeability Fields
-This section evaluates our model’s performance on the baseline dataset 3dp and on a scaling test datapoint; cf.
-Section 3.
-In Section 6.1, we present metrics for isolated velocity (Step 1) and temperature (Step 3) predictions, and for the full pipeline, demonstrating the potential of our approach.
-In ablation studies (Section 6.2), we motivate model design choices.
-Additional metrics and visualizations can be found in Section A.3.
-Table 2: Metrics for predicting $\vec{v}_{\text{pred}}$ in Step 1 and in ablation studies of Step 1 on test and on scaling data from our baseline 3dp dataset with random $k$.
-In [m/y], except for MSE in [m2/y2], SSIM unitless.
-Output Case L MAE MSE SSIM
-LGCNN Step 1
-$v_{x}$ test 190.8046 22.3178 972.5668 0.9911
-$v_{y}$ test 256.2519 32.7444 2031.3357 0.9812
-$v_{x}$ scaling 294.0457 24.9261 1204.1154 0.9911
-$v_{y}$ scaling 367.6891 26.2795 1463.8218 0.9820
-Ablation study Train without data partitioning
-$v_{x}$ test 202.6397 23.8584 1144.4479 0.9913
-$v_{y}$ test 332.9557 35.9249 2490.4104 0.9789
+5 Results
+We evaluate LGCNN’s performance on the baseline dataset 3dp for isolated Steps 1 and 3, and for the full pipeline (end-to-end), demonstrating the potential and scalability of our approach.
+In Section˜5.2, we motivate model design choices through variational experiments.
+Additional metrics and visualizations can be found in Section˜A.3.
+Section˜5.3 explains adaptations, performance, and scalability on the more realistic dataset.
+5.1 Performance on the Synthetic Dataset
+Isolated Steps 1 and 3
+Performance metrics in Table˜1 Ours are for Step 3 (trained and tested on simulated groundtruth $\vec{v}_{\text{sim}}$), and for Step 1 highlighted in gray.
+The combination of quantitative (MAE of (22.3, 32.7) m/y and 0.04 °C, and only 0.07% unconnected cells) and qualitative measures (visual assessment of Figure˜3) demonstrate a strong, physically reasonable performance, especially compared to a generic UNet trained on the same number of datapoints (Figure˜2).
+Figure 3: Step 3: $pki\vec{s}\vec{v}_{\text{sim}}\rightarrow T$ (baseline dataset (3dp): test). 1st column: Label $T$. 2nd-3rd column: LGCNN Step 3, trained without and with zero-padding.
+Full Pipeline
+Testing the full pipeline, i.e., Steps 2 and 3 applied to $\vec{v}_{\text{pred}}$ predicted by Step 1, results in a higher test MAE of 0.09 °C compared to an isolated Step 3.
+This was expected as the isolated Step 3 uses groundtruth, simulated velocity fields $\vec{v}_{\text{sim}}$ as inputs.
+Visual assessment of representative predictions in Figure˜2 reveals physically plausible heat plumes in terms of shape, extent, and heat magnitude, also measurable by a comparably low MoC of 0.09 %.
+Deviations in streamlines arise from smaller errors in the velocity predictions $\vec{v}_{\text{pred}}$ from Step 1, highlighting the input sensitivity of the physical problem, especially near bifurcations, where small perturbations in $\vec{v}$ can lead to an alternative path, e.g., around a clay lens in $k$.
+Compared to training end-to-end on 101dp (see Table˜1 Standard), our model achieves a similar MoC and at least half the accuracy while training on only 1 versus 73 training datapoints, which strongly reduces computation time (see Table˜12 for details) and data requirements.
+Scaling
 […]
-$v_{x}$ test 1047.8464 12.7036 896.2758 0.9016
-$v_{y}$ test 2594.9504 14.3878 2317.6863 0.9903
-$v_{x}$ scaling 659.8439 5.2152 79.2508 0.9699
-$v_{y}$ scaling 456.1904 5.7492 83.9735 0.9948
 […]
-$v_{x}$ test 2453.5881 13.6735 1367.9061 0.9325
-$v_{y}$ test 2029.0269 15.1332 2221.4226 0.9919
-$v_{x}$ scaling 421.4186 8.1184 125.9511 0.9314
-$v_{y}$ scaling 400.2490 7.4200 135.7929 0.9955
-Table 3: Metrics for predicting $T$ in Step 3, in the full pipeline, and in ablations studies of Step 3 on test and on scaling data from our baseline 3dp dataset with random $k$.
-In [°C], except for MSE in [°C2], PAT in [%], SSIM unitless.
-Case L MAE MSE PAT SSIM
-LGCNN Step 3, i.e., test on $\vec{v}_{\text{sim}}$
-test 2.8990 0.0347 0.0041 7.54 0.9304
-scaling 3.0250 0.0168 0.0014 2.05 0.9510
-Pipeline, i.e., test on $\vec{v}_{\text{pred}}$
-test 4.2120 0.0905 0.0307 28.92 0.7637
-scaling 4.9366 0.0413 0.0141 10.87 0.8654
-Ablation study Train in sequence
-test 4.0804 0.0901 0.0289 29.44 0.7553
-Include zero-padding
-test 2.3391 0.0418 0.0048 9.19 0.8838
-Train without data partitioning
-test 3.6420 0.0351 0.0043 7.41 0.9245
+LGCNN obtains a low MAE of 0.02 °C for Step 3, and 0.04 °C for the full pipeline and a very low MoC of 0.06%, see Table˜1.
+A visual assessment of Figure˜4 yields similar qualitative, physically reasonable behavior as the previous results in Figure˜2, demonstrating the scalability of LGCNN.
+5.2 Experiments on the Synthetic Dataset
+Training in Sequence
+When training Step 3, we consider two alternatives for the velocity input: simulated velocities $\vec{v}_{\text{sim}}$ and velocities predicted by Step 1 $\vec{v}_{\text{pred}}$.
+Training the model sequentially using $\vec{v}_{\text{pred}}$ results in substantially higher MoC values compared to training with $\vec{v}_{\text{sim}}$, as shown in Table˜1 (Experiments vs.
+Ours).
+As illustrated in Figure˜2, using $\vec{v}_{\text{pred}}$ during training leads to physically implausible temperature fields characterized by noisy artifacts and fragmented heat plumes.
+We attribute this behavior to local misalignments between the predicted streamline inputs $\vec{s}$ and the corresponding temperature label $T$, which strongly complicates localized learning.
+In contrast, training with $\vec{v}_{\text{sim}}$ produces streamline fields that are physically consistent with the temperature distribution which provides informative gradients and enables stable and effective training.
+Enabling Zero-Padding
+Using no padding in all convolutional layers produces smaller but cleaner output fields, as shown in Figure˜3.
+This leads to a slight improvement in standard ML metrics and a substantial improvement in the MoC score (Table˜1, Ours vs.
+Experiments), indicating fewer disconnected heat plumes.
+Partitioning Training Data
+Hyperparameter search shows the best performance for partitioning the domain during training into 20 736 overlapping patches of $\times256$ cells for Step 1 and 82 944 patches for Step 3.
+Although this increases per-epoch compute time compared to full-domain training, it significantly reduces the number of epochs needed to reach comparable test performance (cf.
+Table˜1 Experiments vs.
+Ours, Step 1), resulting in an overall training time reduction of 67–90% (cf.
+Table˜11) while maintaining comparable accuracy across all metrics.
 […]
-test 4.4658 0.0416 0.0067 11.69 0.9835
-scaling 4.3489 0.0176 0.0022 3.07 0.9954
-[…]
-test 4.2772 0.0376 0.0058 10.96 0.9915
-scaling 3.9946 0.0192 0.0025 4.73 0.9968
-6.1 Performance of LGCNN
-Isolated Steps 1 and 3
-Performance metrics for Step 1 are shown in Table 3, for Step 3 (trained and tested on simulated $\vec{v}_{\text{sim}}$) in Table 3.
-The combination of quantitative (MAE of (22.3, 32.7) m/y and 0.03 °C) and qualitative results through a visual assessment of Figure 3 demonstrate a strong performance with heat plumes forming according to $k$.
-Figure 3: $pki\vec{s}\vec{v}_{\text{sim}}\rightarrow T$ (baseline dataset (3dp), test). 1st column: Label $T$. 2nd-3rd column: LGCNN Step 3, trained without and with zero-padding.
-Full Pipeline
-Testing the full pipeline, i.e., Steps 2 and 3 are based on $\vec{v}_{\text{pred}}$, results in a higher test MAE of 0.09 °C compared to an isolated Step 3 (see Table 3).
-This was expected as the isolated Step 3 uses the true, simulated velocity fields $\vec{v}_{\text{sim}}$ as inputs.
-Visual assessment of representative predictions in Figure 2 reveals physically plausible heat plumes in terms of shape, extent, and heat magnitude.
-Deviations in streamlines arise from smaller errors in the velocity predictions from Step 1 $\vec{v}_{\text{pred}}$, highlighting the input sensitivity, especially near bifurcations, where small perturbations in $\vec{v}$ can lead to an alternative path around a clay lens in $k$.
-Compared to fully data-driven models trained on 101dp (see Table 1), our model achieves a similar accuracy while training on only 1 versus 73 training datapoints, which strongly reduces computation time (see Table 16 for details).
-Scaling Test
-We can already expect that our model is able to scale to larger domains, as we train only on local patches of the spatial domain and achieve high test accuracies on the full domain during validation and testing; see also Section 6.2 Training Data Partitioning for more details.
-To further demonstrate the scaling, we test our model on a domain of 4$\times$ the size of the training domain with the same number of heat pumps, hence with a lower density of GWHPs/km2; cf.
-Figure 4.
-[…]
-We obtain a comparable MAE of 0.02 °C for Step 3, and an MAE of 0.04 °C for the full pipeline, see Table 3.
-A visual assessment of Figure 4 yields similar qualitative behavior as the previous results in Figure 2.
-6.2 Ablation Study
-Training in Sequence
-[…]
-Zero-Padding
-Removing zero-padding from all convolutional layers yields smaller but cleaner output fields, as shown in Figure 3.
-As a result, we obtain improved performance, reflected in lower PAT and higher SSIM scores in Table 3.
-Partitioning Training Data
-Hyperparameter search shows the best performance for partitioning the domain during training into 20 736 overlapping patches of 256$\times$256 cells for Step 1 and 82 944 patches for Step 3.
-Although this increases per-epoch compute time compared to full-domain training, it significantly reduces the number of epochs needed to reach comparable test performance (cf.
-Tables 3 and 3), resulting in an overall training time reduction of 67–90% (cf.
-Table 15).
-[…]
-We individually replace each of the three steps with an optimized (cf.
-Section A.2.1) vanilla approach, i.e., UNet or DDUNet trained on 101dp.
-For Step 2, both fail completely; results for Steps 1 and 3 are summarized in Tables 3 and 3.
-In Step 1, UNet slightly outperforms DDUNet, while the reverse holds for Step 3.
-A visual comparison of the scaling datapoint in Figure 4 reveals that increasing training data from 3dp to 101dp has minimal impact on the performance in Step 3.
+We replace each step with an optimized UNet or DDUNet trained on 101dp (cf.
+Section˜A.2.3).
+Step 2 fails for both models.
+Results for Steps 1 and 3 are summarized in Table˜1 Experiments.
+In Step 1, both models struggle to predict $v_{y}$ due to outliers in training data.
+In Step 3, performance is comparable to 3dp in standard metrics but yields a substantially higher MoC, with scaling results in Figure˜4 showing minimal benefit from increased training data.
 Modifying Streamline Inputs of Step 3
@@
 Omitting only $s_{o}$ results in overly narrow plumes, whereas excluding only $s$ produces overly smeared temperature fields.
-Including both but removing the fading of values along the streamlines leads to diffuse background temperatures and less localized plumes.
-Overall, the Huber validation loss increases by 32–132%, depending on the experiment.
-For visual, quantitative, and extended results, we refer to Section A.3.
-7 Domain Transfer to Real Permeability Fields
-Finally, we investigate whether our approach transfers to the realistic dataset in Section 3, which is based on real-world permeability fields $k$ that are cut from geological maps of Munich.
-Since these maps cover a limited area, the number of potential datapoints is very small.
-We compare performance for temperature prediction of Step 3 and the full pipeline on the validation and scaling datapoints.
-Results are presented in Tables 5 and 5; additional metrics are in Section A.3.
-Table 4: Metrics for predicting $\vec{v}_{\text{pred}}$ in Step 1 on validation and scaling data from our realistic dataset. In [m/y], except for MSE in [m2/y2], SSIM unitless.
-Output Case L MAE MSE SSIM
-Step 1
-$v_{x}$ val 106.8607 15.4095 380.3762 0.9939
-$v_{y}$ val 74.5570 10.6605 148.2475 0.9993
-$v_{x}$ scaling 459.1620 110.0078 13570.0000 0.9462
-$v_{y}$ scaling 240.9406 17.6051 616.5186 0.9965
-Table 5: Metrics for predicting $T$ in Step 3 and in the full pipeline on validation and scaling data from our realistic dataset. In [°C], except for MSE in [°C2], PAT in [%], SSIM unitless.
-Case L MAE MSE PAT SSIM
-Step 3, i.e., test on $\vec{v}_{\text{sim}}$
-val 0.8222 0.0175 1.01e-3 2.18 0.9672
-scaling 0.8052 0.0189 8.21e-4 0.92 0.9497
-Full pipeline, i.e., test on $\vec{v}_{\text{pred}}$
-val 2.3194 0.0841 2.75e-2 27.79 0.7510
-scaling 2.0511 0.0394 4.42e-3 10.02 0.8708
+Including both but removing the fading over time along the streamlines leads to diffuse background temperatures and overly precise and too long plumes.
+Overall, the Huber validation loss increases by 32-132%, depending on the experiment.
+For extended and visual results, we refer to Section˜A.3.
+5.3 Domain Transfer to Real Permeability Fields
+We evaluate our method on real permeability fields $k$ (Section˜2), for which only four datapoints are available.
+Results for Step 3 and the full pipeline on the validation and scaling datapoints are shown in Table˜1 Domain transfer; additional metrics are provided in Section˜A.3.
 Methodological Adaptations
-Slight adjustments to chunk size, overlap, data split, and architecture yield better performance compared to those used on the baseline dataset, see Section A.2.1 for details.
-Visual inspection of input $k$ in Figure 5 reveals fewer but larger-scale features.
+Slight adjustments to chunk size, overlap, data split, and architecture yield better performance compared to those used on the baseline dataset, see Section˜A.2.3 for details.
+Visual inspection of input $k$ in Figure˜5 reveals fewer but larger-scale features.
 Therefore, training can benefit from a larger spatial context, which requires more training data.
 We increase the number of training samples by using three out of the four available datapoints for training and one for validation.
-For the streamlines computation, we switch to an explicit fourth-order Runge–Kutta scheme, which proved stable and sufficient for training and inference due to the lower frequency in $k$.
+For the streamlines computation, we switch to an explicit fourth-order Runge–Kutta scheme, which proved stable and sufficient for training and inference due to the lower frequencies in $k$.
 Performance and Scaling Tests
-In Step 3, we reach an even lower maximum absolute error $L_{\infty}$ of 0.82 °C compared to 2.90 °C of our initial model on the baseline dataset (Section 6).
-Other metrics confirm a similar performance; see Table 3.
-In the scaling test, most errors are comparable too.
-For the full pipeline, errors are slightly higher on the validation and scaling data compared to an isolated Step 3, consistent with our observations on the baseline dataset (Table 3).
-Physical consistency, which is our primary objective, remains strong: Shape, magnitude, and connectivity of the predicted heat plumes are well preserved, even in the scaling test (Figure 5).
+In Step 3, we reach an even lower maximum absolute error $L_{\infty}$ of 0.82 °C compared to 2.90 °C of our initial model on the baseline dataset (Section˜5).
+Other metrics confirm this except for SSIM and MoC; cf.
+Table˜1 Domain transfer.
+In the scaling test, most errors behave as before, as expected.
+For the full pipeline, errors are slightly higher on the validation and scaling data compared to an isolated Step 3, consistent with our observations on the baseline dataset.
+Physical consistency, which is our primary objective, remains strong: Shape, magnitude, and connectivity of the predicted heat plumes are well preserved, even in the scaling test (Figure˜5).
 Local deviations are primarily due to differences between $\vec{v}_{\text{sim}}$ and $\vec{v}_{\text{pred}}$, as evident from the discrepancy between Step 3 and full pipeline outputs.
-Figure 5: $pki\rightarrow T$ (realistic dataset, scaling). 1st & 2nd column: Label $T$ and input $k$. 3rd-4th column: LGCNN (isolated Step 3 and full pipeline).
+Figure 5: Predictions of $T$ (realistic dataset (4dp): scaling). Input $k$, label $T$ LGCNN isolated Step 3, and full pipeline.
+Based on their poor generalization in data-scarce regimes observed on the baseline dataset, we omit generic ML architectures from the experiments on realistic $k$ fields.
 6 Conclusions and Future Work
-In real-world scientific applications, data is often very limited.
-At the same time, models are expected to generalize and scale to large domain sizes, pushing standard data-driven models to their limits.
-To address this issue, we proposed the LGCNN model, a hybrid, physics-inspired CNN approach.
-In the prediction of groundwater temperature fields under arbitrary heat pump configurations and real permeability data, we identify long-range interactions driven by advection as the problematic process for classical deep learning models, such as vanilla CNNs.
-We replace this part of the prediction with a simple numerical solver, resulting in a sequence (CNN 1 – numerical solver – CNN 2).
-While this setup is specific to groundwater heat transport, the general approach should be applied to a wider range of applications.
-Finally, we stress that traditional error metrics can obscure effects like input sensitivity; therefore, we additionally rely on visual assessments and problem-specific metrics such as PAT.
-In future work, we will address dataset limitations, investigate input sensitivity, and accelerate streamline computation.
+We introduce the Local-Global Convolutional Neural Network (LGCNN) to address complex, data-scarce learning problems governed by both local and non-local physical interactions.
+Trained on only a few simulations, LGCNN accurately predicts subsurface temperature fields on much larger two-dimensional heterogeneous domains, while providing an inference-time speed-up of roughly 2,000× compared to classical high-fidelity simulations on the training data domain.
+By combining translation-invariant CNNs for local processes with a lightweight numerical surrogate for global transport, our method achieves data-efficient, scalable, and physically consistent predictions, outperforming conventional ML models such as UNet variants that struggle to capture long-range dependencies in data-scarce regimes.
+Although demonstrated for groundwater heat transport, the concept of local-global decomposition can be generalized to a broad class of advection-diffusion systems.
+Future work will extend the approach to three-dimensional and transient scenarios, investigate sensitivity to input uncertainty, and further accelerate streamline computation.
 Acknowledgements
@@
 We acknowledge the support by the Stuttgart Center for Simulation Science (SimTech).
-We also acknowledge the use of computational resources of the DelftBlue supercomputer, provided by Delft High Performance Computing Centre [DHPC2024].
+We also acknowledge the use of computational resources of the DelftBlue supercomputer, provided by Delft High Performance Computing Centre (Delft High Performance Computing Centre, DHPC).
 We thank Kai Zosseder and Fabian Böttcher at the Technical University of Munich for their application-related input.
+Reproducibility
+To ensure reproducibility, our code to train and evaluate models, the trained and evaluated models themselves, and the datasets designed for training and testing are included in the supplementary material and will be published after the review process.
+Our raw data and pretrained models are publicly available on SurfDrive, see below in Supplementary Material.
+Our code for data generation (https://github.com/JuliaPelzer/Dataset-generation-with-Pflotran) and model training (including dataset preparation and evaluation routines) (for LGCNN and data preparation: https://github.com/JuliaPelzer/Heat-Plume-Prediction/tree/AllIn1/LGCNN/release25, for DDUNet: https://github.com/corne00/DDUNetForHeatplumePrediction) is on GitHub.
+Follow the instructions in the respective readme files to prepare the data and train or infer a model.
+Supplementary Material
 […]
-Data and Code Accessibility
-Our raw data and pretrained models are publicly available on DaRUS https://darus.uni-stuttgart.de/dataverse/ipvs_lgcnn/ or more specifically:
 Raw datasets:
-Pretrained models (including hyperparameters):
+  • Dataset of random permeability with 3+1 datapoints (3dp + 1)
+  • Dataset of random permeability with 101 datapoints (101dp)
+  • Dataset of real permeability with 4+1 datapoints (4dp + 1)
+  • Prepared dataset for training the FNO, because the other datapoints are too large
+Trained models (including hyperparameters):
 […]
 […]
-  • LGCNN on real permeability fields, 4dp: https://doi.org/10.18419/darus-5082
-Our code for data generation (https://github.com/JuliaPelzer/Dataset-generation-with-Pflotran) and model training (including dataset preparation and evaluation routines) (for LGCNN and data preparation: https://github.com/JuliaPelzer/Heat-Plume-Prediction/tree/AllIn1/LGCNN/release25, for DDUNet: https://github.com/corne00/DDUNetForHeatplumePrediction) is on GitHub.
-Follow the instructions in the respective readme files to prepare the data and train or infer a model.
+  • LGCNN on random permeability fields, 3dp
+[…]
+  • LGCNN on real permeability fields, 4dp
+  • vanilla models first trying to overfit them on 1 data point: PI-CNN, FNO
+Code (including training and evaluation routines):
+[…]
+[…]
+[…]
+All supplementary material can be accessed via SURFDrive
+under this link: https://surfdrive.surf.nl/files/index.php/s/f3Oqg3ufir9T9LL.
 Appendix A Appendix
 A.1 Simulation Setup and Transport Regime
-In Section 3, we describe how the two datasets are generated using the subsurface simulation software Pflotran [pflotran], which solves the coupled mass and energy conservation equations.
+In Section˜2, we describe how the two datasets are generated using the subsurface simulation software Pflotran (Lichtner et al., 2015a), which solves the coupled mass and energy conservation equations.
 Here, we provide additional technical details and modeling assumptions relevant for reproducibility.
@@
 Hydro-geological, Operational and Simulation Parameters
-Our data was generated on 12.8km×$\times$12.8km×$\times$5m with a cell size of 5m×$\times$5m×$\times$5m for the baseline simulations.
+Our data was generated on $\times12.8km$$\times$5m with a cell size of $\times5m$$\times$5m for the baseline simulations.
 To test scalability, we also simulate a domain that is twice as large in both $x$- and $y$-dimension for the baseline dataset, but only twice as large in $y$-dimension for the more realistic dataset due to dimension restrictions of the available input data.
-The baseline dataset uses a constant hydraulic pressure gradient $\nabla p$ of 0.003 [pelzer2024, geologica]; permeability field $k$ is generated using Perlin noise [perlin] within (1.02E-11, 5.10E-9) m2, and $100$ randomly placed heat pumps, which are all operating with a 5 i.e., C injection temperature difference compared to the surrounding groundwater and an injection rate of 0.00024 m3/s.
-All values are based on realistic parameter ranges in the region of Munich[geologica, geokw22].
-For the realistic dataset, instead of random permeability fields $k$, we use permeability field data that is derived from borehole measurements in the Munich region[lfu_gepo].
+The baseline dataset uses a constant hydraulic pressure gradient $\nabla p$ of 0.003 (Pelzer & Schulte, 2024; geo, 2022); permeability field $k$ is generated using Perlin noise (Perlin, 1985) within (1.02E-11, 5.10E-9) m2, and $100$ randomly placed heat pumps, which are all operating with a 5 °C injection temperature difference compared to the surrounding groundwater and an injection rate of 0.00024 m3/s.
+All values are based on realistic parameter ranges in the region of Munich(geo, 2022; Zosseder et al., 2022).
+For the realistic dataset, instead of random permeability fields $k$, we use permeability field data that is derived from borehole measurements in the Munich region(Bayerisches Landesamt für Umwelt, 2015).
 Furthermore, based on subsurface measurements, we set $\nabla p=0.0025$ for these simulations.
 All other simulation parameters remain identical to the baseline.
-Other subsurface and operational parameters are taken directly from [pelzer2024], which also includes additional information about solver setup and boundaries.
-For mathematical details of the governing equations, we refer the reader to [pflotran_theory, anderson2007introducing, delleur2016elementary].
-Simplifications
+Other subsurface and operational parameters are taken directly from Pelzer & Schulte (2024), which also includes additional information about solver setup and boundaries.
+For mathematical details of the governing equations, we refer the reader to Lichtner et al.
+(2015b); Anderson (2007); Delleur (2016).
+Notation Simplification
 For simplifications, we refer to the (hydraulic) pressure field $p$ throughout the paper.
@@
 Péclet Number
-To quantify whether heat transport in our system is dominated by advection or diffusion, we compute the dimensionless Péclet number [rapp2017peclet], which is defined as
+To quantify whether heat transport in our system is dominated by advection or diffusion, we compute the dimensionless Péclet number (Rapp, 2017), which is defined as
 $\text{Pe}=\frac{L\cdot v}{\alpha},$
 with the characteristic length $L$, the local velocity $v$, and the thermal diffusivity $\alpha$, defined as
-$\alpha=\frac{k}{\rho c_{p}},$
-with $k$ the thermal conductivity, $\rho$ the density, and $c_{p}$ the specific heat capacity. We take the parameters of the solid phase of our simulation inputs of
-  • Thermal conductivity: $k=0.65\,\mathrm{W/(m\cdot K)}$
+$\alpha=\frac{\kappa}{\rho c_{p}},$
+with $\kappa$ the thermal conductivity, $\rho$ the density, and $c_{p}$ the specific heat capacity. We take the parameters of the solid phase of our simulation inputs of
+  • Thermal conductivity: $\kappa=0.65\,\mathrm{W/(m\cdot K)}$
   • Density: $\rho=2800\,\mathrm{kg/m^{3}}$
@@
   • Mean: $330\,\mathrm{m/year}\approx 1.04\times 10^{-5}\,\mathrm{m/s}$
-[…]
+With a characteristic length determined by the heat plume length of $2000\text{--}6000\,\mathrm{m}$ and mean x-velocity, we get:
 $Pe=\frac{L\cdot v}{\alpha}\approx L\cdot 10^{2}\ \approx 10^{5}$
@@
 Therefore, $\text{Pe}\approx 10^{5}$ indicates that in the simulations, the heat transport is advection-dominated at larger scales.
-A.2 Neural Networks, Hyperparameters and Training Details
-This section provides implementation and training details for all neural network models used in this work.
+Reasoning for Decoupling Assumptions
+In LGCNN, the goal is to split properties into local and global.
+To build separate datasets for the isolated steps, we require reasonable inputs and labels.
+Hence, the from data or simulation generated physical properties are analyzed as potential interim inputs / labels.
+From simulation, we get $\vec{v}$ (mostly local) and $T$ (heat is transported in an advective and diffusive manner $\rightarrow$ local and global).
+To split the temperature field into local and global, we handcraft a new property streamlines $\vec{s}$ that describe the central evolution of a heat plume driven only by advection, starting at each heat pump location.
+An important assumption underlying our approach is that the coupling between velocity and temperature can predominantly be treated as one-way in our setting.
+The velocity field strongly influences the temperature distribution but the temperature also influences the velocity.
+Since this effect mostly occurs in 3D domains with very high temperature gradients and we model 2D and very small temperature gradients, our assumption holds.
+The general background velocity field influences the streamlines directly but the streamlines or rather the injection points and their mass influx also disturbs the local velocity field which in turn can influence other streamlines passing by from further upstream.
+By including the start locations of the streamlines, i.e., the locations of the heat pumps, as inputs to $\vec{v}$ and taking the steady-state flow field with running heat pumps from the simulation as a label, the influence of $\vec{s}$ is reduced and again decoupled as an additional input $i$ with the influence only local and already included in the labels.
+Hence, a one-way coupling from positions of heat pumps $i\rightarrow\vec{v}\rightarrow\vec{s}$ is possible.
+A.2 NNs, Hyperparameters and Training Details
+This section provides implementation and training details for all NN models used in this work.
 We begin with a description of the baseline UNet architecture, which also forms the foundation of both LGCNN and DDU-Net.
 Then, we outline the hyperparameter optimization process and present the final configurations per model and dataset.
-All models are trained using PyTorch[paszke2017automatic].
-UNet Architecture
+All models are trained using PyTorch (Paszke et al., 2017).
+A.2.1 UNet Architecture
 The UNet architecture used in this work is based on the original design presented by Ronneberger et al.
-[ronneberger2015u], with several modifications to tailor it to the specific needs of our task of predicting temperature fields.
+(2015b), with several modifications to tailor it to the specific needs of our task of predicting temperature fields.
 In this section, we introduce the key concepts that define the architecture and explain how they contribute to the model’s design.
@@
 The weight decay parameter helps prevent overfitting by penalizing large weights in the model.
-Furthermore, the Adam optimizer [kingma2014adam] is employed as optimizer.
-For the realistic permeability field, we additionally introduce the optimizable hyperparameter optimizer switch, which, when enabled, switches the optimizer to LBFGS after 90 epochs.
+Furthermore, the Adam optimizer (Kingma, 2014) is employed as optimizer.
+For the realistic permeability field, we additionally introduce the optimizable hyperparameter optimizer switch, which, when enabled, switches the optimizer to limited-memory BFGS (LBFGS) after 90 epochs.
 During inference, each datapoint is processed as a whole, but during training, they are loaded in smaller patches cut out from the datapoint.
-Optimized hyperparameters include their overlap, i.e., inverse of skip per direction, and their size, called box length.
-The data split is untouched by this, i.e.
-one simulation run per train / val / test separately.
+Optimized hyperparameters include their overlap, i.e., inverse of skip per direction, and the patch size, called box length.
+The data split is untouched by this, i.e., separate simulation runs for train / val / test.
 This is important to not overlap test patches with training ones.
-A.2.1 Hyperparameters and Hyperparameter Optimization
-We optimize the hyperparameters of our architecture, training process and data loading in several rounds with Optuna [optuna_2019] and additional manual tweaking.
+A.2.2 Energy-Based Loss for PINNs
+For our PINN, we overfitted a UNet with an energy loss $\mathcal{L}_{\text{energy}}$ to a single datapoint from 3dp.
+The physics-informed loss for predicting the temperature field $T$ considers energy conservation, including advective and conductive terms, and heat pump contributions.
+Let $\vec{v}=(v_{x},v_{y})$ be the velocity field and $p$ the pressure.
+Denote the predicted temperature as $\hat{T}$.
+The energy loss is defined as
+$\mathcal{L}_{\text{energy}}=\text{MSE}\Big(E-E^{\text{HP}}_{\text{conv}},\,0\Big),$ (5)
+where $E$ is the local energy residual and $E^{\text{HP}}$ the energy inflow from heat pumps.
+Energy residual.
+The energy residual accounts for advection and thermal diffusion:
+$\rho_{m}$ $=\text{molar density}(\hat{T},p),$ (6)
+$h$ $=\text{enthalpy}(\hat{T},p),$ (7)
+$E_{u}$ $=\frac{\partial}{\partial x}\Big(\rho_{m}\,h\,v_{x}-\frac{\kappa}{\Delta x}\frac{\partial\hat{T}}{\partial x}\Big),$ (8)
+$E_{v}$ $=\frac{\partial}{\partial y}\Big(\rho_{m}\,h\,v_{y}-\frac{\kappa}{\Delta x}\frac{\partial\hat{T}}{\partial y}\Big),$ (9)
+$E$ $=E_{u}+E_{v},$ (10)
+where $\kappa$ is the thermal conductivity and $\Delta x$ is the uniform grid spacing.
+Heat pump contribution.
+Heat pumps add localized energy sources based on a one-hot-encoding of the positions of heat pumps (ID) to get a spatially distributed contribution:
+$\rho$ $=\text{density}(\hat{T},p),$ (11)
+$E^{\text{HP}}$ $=c_{p}\,\rho\,\Delta T\,Q\,\frac{\text{ID}}{\Delta x^{3}},$ (12)
+$E^{\text{HP}}_{\text{conv}}$ $=K*E^{\text{HP}},$ (13)
+where $c_{p}$ is the specific heat capacity of water, $\Delta T$ is the temperature difference at the injection well, $Q$ is the volumetric flow rate, ID is the material indicator (1 if heat pump present, 0 otherwise), and $K$ is a fixed, handcrafted $3\times 3$-convolutional kernel $K=\begin{bmatrix}-1&0&1\\ 0&0&0\\ 1&0&-1\end{bmatrix}$ with zero-padding used to smooth the inflow contribution.
+It distributes the inflow over neighboring cells while preserving directional structure and its shape matches our observations.
+A.2.3 Hyperparameters and Hyperparameter Optimization
+We optimize the hyperparameters of our architecture, training process and data loading in several rounds with an automated tree-based search using Optuna (Akiba et al., 2019) and additional manual tweaking.
 Optuna performs optimization using a tree-structured Parzen estimator algorithm.
@@
 […]
-The inputs to the model are $pki$ (pressure field, permeability field, and location of heat pumps).
-For training, the Mean Absolute Error (MAE) loss function was used, and optimized with the Adam optimizer.
+The inputs to the model are pkipki (pressure field, permeability field, and location of heat pumps).
+For training, the MAE loss function was used, and optimized with the Adam optimizer.
 The learning rate is set to $1\times 10^{-5}$.
 The model is trained for 10,000 epochs.
-Hyperparameters for purely data-driven approaches
-The values considered during hyperparameter optimization for the UNet and DDUNet, trained on 73 datapoints and performed using Optuna [optuna_2019], along with the best settings found, are listed in Table 6.
+Hyperparameters for UNet variants on 101dp
+The values considered during hyperparameter optimization for the UNet and DDUNet, trained on 73 datapoints and performed using Optuna (Akiba et al., 2019), along with the best settings found, are listed in Table˜2.
 Certain hyperparameters were fixed: we set the number of epochs to 750, with an early stopping criterion based on validation loss and a patience of 80 epochs.
@@
 Note that the number of communicated feature maps in the vanilla UNet is simply an extra convolution layer in the coarsest part of the UNet (without communication).
-Hyperparameter Range $pki\rightarrow v_{x}v_{y}$    (Step 1) $kiv_{x}v_{y}ss_{o}\rightarrow T$    (Step 3) $pki\rightarrow T$    (Full)
+Hyperparameter Range $pki\rightarrow v_{x}v_{y}$ (Step 1) $iv_{x}v_{y}ss_{o}k\rightarrow T$ (Step 3) $pki\rightarrow T$ (Full)
 UNet101dp $2\times 2$}DDUNet101dp UNet101dp $2\times 2$}DDUNet101dp UNet101dp $2\times 2$}DDUNet101dp
@@
 After the hyperparameter search, the values corresponding to the best-performing model (based on Huber loss for the validation dataset) were selected.
-With these values fixed, five models were trained using different randomly sampled initializations to evaluate sensitivity to random initialization, for these values see Tables 11 and 10.
-LGCNN Hyperparameters - on Random Permeability
-The values considered during hyperparameter optimization with Optuna [optuna_2019] and the best settings found for both steps of LGCNN are listed in Table 7.
+With these values fixed, five models were trained using different randomly sampled initializations to evaluate sensitivity to random initialization, for these values see Tables˜8 and 7.
+Hyperparameters for UNet+++ and dilated UNet
+[…]
+Table 3: Baseline hyperparameters for end-to-end temperature prediction.
+Parameter Dilated CNN (3dp) UNet++ (101dp)
+Parameters of the Dataset
+dataset size 3 101
+inputs pikpik ikik
+cutouts / patches full domain full domain
+Parameters of Training
+batch size 20 2
+training loss MAE MSE
+initial learning rate $10^{-4}$ (manual decay) $2\times 10^{-4}$
+optimizer Adam Adam
+epochs 1000 190
+Parameters of the Network
+architecture Dilated CNN (zero-padded) UNet++
+kernel size 5 default
+depth 3 default
+initial features 32 default
+stride 1 default
+dilation 10 default
+activation function ReLU default
+normalization BatchNorm default
+repeat inner blocks False -
+Hyperparameters for PINN and FNO
+Hyperparameters of our PI-CNN and FNO trainings are listed in Table˜4.
+For PI-CNN, we used the optimized, well-working architecture from before; for FNO a transferred set of hyperparameters with an additional hyperparameter search on layers, modes, inputs and optimizer.
+Table 4: Baseline hyperparameters for overfitting experiments on a single datapoint (dp id 0).
+Parameter FNO (overfit test) PINN (overfit test)
+Parameters of the Dataset
+dataset size 1 1
+inputs pikpik $pikv_{x}v_{y}$
+box length 480 256
+Parameters of Training
+batch size 8 20
+training loss MSE MSE (physics-only)
+initial learning rate $10^{-3}$ (LBFGS) $10^{-4}$ (manual decay)
+optimizer LBFGS Adam
+epochs 1000 500
+Parameters of the Network
+architecture FNO UNet (zero-padding)
+modes / kernel size 64 modes $4\times 4$
+depth 6 4
+initial features 16 32
+stride 1 1
+dilation 1 1
+activation function ReLU ReLU
+normalization BatchNorm BatchNorm
+repeat inner blocks False False
+Hyperparameters for LGCNN on random permeability
+The values considered during hyperparameter optimization with Optuna (Akiba et al., 2019) and the best settings found for both steps of LGCNN are listed in Table˜5.
 Although the optimization was originally run for 100 epochs, the optimum was consistently found within the first 25 epochs.
@@
 Table 5: LGCNN-Random $k$: Hyperparameter optimization: Parameter ranges and best configurations.
-Parameter Range 1st step 3rd step
+Parameter Range Step 1 ($v$) Step 3 ($T$)
 Parameters of the Dataset
-inputs $v:$ $(p,i,k)$ $pik$ $iv_{x}v_{y}ss_{o}k$
+inputs $v:$ $(p,i,k)$ pikpik $iv_{x}v_{y}ss_{o}k$
 $T:$ $(i,v_{x},v_{y},s,s_{o},k)$
@@
 kernel size 3, 4, 5 5 4
-depth $v:$ 1, 2, 3, 4 4 4
-$T:$ 1, 2, 3
-Fixed parameters for this hyperparameter search are the learning rate (fixed at $10^{-4}$), ReLU as activation function, the batch size of 20, and the use of a batch normalization layer within the inner blocks of the UNet architeture.
+depth 1, 2, 3, 4 4 4
+Fixed parameters for this hyperparameter search are the learning rate (fixed at $10^{-4}$), ReLU as activation function, the batch size of 20, and the use of a batch normalization layer within the inner blocks of the UNet architecture.
 The validation loss used for selecting the optimal model is the MAE.
-LGCNN Hyperparameters - on Real Permeability
-The values considered during hyperparameter optimization on the dataset with a more realistic permeability field were selected using Optuna [optuna_2019], and are summarized in  Table 8, along with the best configurations found for both steps of the LGCNN.
+Hyperparameters for LGCNN on real permeability fields
+The values considered during hyperparameter optimization on the dataset with a more realistic permeability field were selected using Optuna (Akiba et al., 2019), and are summarized in  Table˜6, along with the best configurations found for both steps of the LGCNN.
 The optimization was run for up to 100 epochs.
 For more background on the network architecture and the various hyperparameters, cf.
-A.2.
+A.2.1.
 Several hyperparameters were fixed during this process.
@@
 Parameters of the Dataset
-inputs $v:$ $ki,pik,gik,gk,pk$ $pik$
+inputs $v:$ $\nabla py,p,k,i$ pikpik
 $T:$ $iv_{x}v_{y}ss_{o}k$ $iv_{x}v_{y}ss_{o}k$
@@
 optimizer switch True, False False False
-learning rate 1e-3, 5e-4, 1e-4, 5e-5 1e-3 1e-4
+learning rate 1e-3, 5e-4, 1e-4, 5e-5 1e-4 1e-4
 Parameters of the Network
-No. initial features 8, 16, 32 32 16
-kernel size 3, 5 3 5
-depth 4, 5, 6 5 6
+No. initial features 8, 16, 32 16 16
+kernel size 3, 5 5 5
+depth 4, 5, 6 6 6
 repeat inner True, False False False
@@
 This section provides additional experimental results.
-While the main results section focused only on the test and scaling datasets, we also include here the metric values on the training and validation datasets.
-For completeness and easier comparison, the test and scaling metrics are re-listed as well.
-Purely data-driven UNet, DDU-Net: Metrics of training and ablation study
-Table 9 presents all the metrics for predicting the temperature field directly from the inputs $pki$ using a data-driven approach, evaluated on the training, validation, and test datasets.
-The results are provided for several models: (1) UNet trained on only 1 datapoint and tested and validated on 2 additional datapoints, (2) a UNet trained on 73 datapoints (73-18-10 train-validation-test split), and (3) a DDUNet trained on the same 73 datapoints dataset, operating on $2\times 2$ subdomains.
-Furthermore, to assess the model’s sensitivity to random initialization, the training of the same architecture was repeated five times for the most relevant models.
+While the main results section focused only on the test and scaling datasets, we also include here the metric values on training and validation datasets.
+For easier comparison, the test and scaling metrics are re-listed as well.
+Vanilla UNet and DDUNet on 101dp: Metrics of training and ablation study
+We present metrics for predicting the temperature field directly from the inputs pkipki using a data-driven approach, evaluated on the training, validation, and test datasets.
+The results are provided for two models: (1) UNet trained on 73 datapoints (73-18-10 train-validation-test split), and (2) DDUNet trained on the same 73 datapoints dataset, operating on $2\times 2$ subdomains.
+To assess the model’s sensitivity to random initialization, the training of the same architecture was repeated five times.
 Based on these repetitions, the mean and standard deviation of the performance metrics were computed using the following equations:
@@
 where $x_{i}$ denotes the metric value from the $i$-th training run, and $n=5$ is the number of runs.
-These results are summarized in Table 10 and Table 11.
+These results are summarized in Table˜7.
 The choice of $n=5$ was made empirically to balance computational effort and statistical reliability.
-The standard deviations in  Tables 11 and 10 were used as validation: they are neither excessively large (indicating instability) nor unrealistically small (indicating insufficient sampling).
-In addition to testing the UNet and DDUNet trained on 73 datapoints on the 10-point test dataset, we also evaluate these models on the same datapoint used to test the UNet3DP.
-Table 9: Performance metrics for predicting $T$ with different models and datasets.
-Errors in [C], MSE in [C], PATs in [%] and SSIM unitless.
-The LGCNN-test dataset corresponds to the 1 datapoint used for testing the LGCNN approach.
-Model Data Case Huber L MAE MSE PAT SSIM
-UNet3DP randomK3 train 0.0020 2.4954 0.0404 0.0040 6.43 0.8281
-val 0.0269 5.2901 0.1365 0.0574 38.82 0.5717
-test 0.0235 4.8642 0.1314 0.0492 39.05 0.5794
-UNet rK101 train 0.0010 4.2443 0.0172 0.0021 2.34 0.9960
-val 0.0051 4.1972 0.0441 0.0106 12.68 0.9859
-test 0.0050 4.2140 0.0426 0.0104 11.90 0.9869
-* LGCNN-test 0.0048 4.3985 0.0473 0.0100 13.63 0.9827
-* scaling 0.0016 4.3426 0.0202 0.0033 4.33 0.9955
-2$\times$2-DDUNet rK101 train 0.0018 3.8100 0.0236 0.0038 4.47 0.9940
-val 0.0079 4.0084 0.0550 0.0165 17.03 0.9825
-test 0.0076 3.7006 0.0549 0.0159 17.14 0.9835
-* LGCNN-test 0.0063 3.4257 0.0548 0.0128 17.42 0.9804
-* scaling 0.0025 4.1806 0.0235 0.0052 6.11 0.9940
-Table 10: Statistics for predicting $T$ with different models and datasets.
-Errors in [C], MSE in [C], PATs in [%] and SSIM unitless.
+The standard deviations in  Table˜7 were used as validation: they are neither excessively large (indicating instability) nor unrealistically small (indicating insufficient sampling).
+In addition to testing the UNet and DDUNet trained on 73 datapoints on the 101dp test dataset, we also evaluate these models on the same datapoint used to test the UNet3dp.
+Table 7: Statistical error metrics for end-to-end prediction of $T$ with 101dp UNet and DDUNet.
+Errors in [°C], MSE in [°C2], and SSIM unitless.
 The LGCNN-test dataset corresponds to the 1 datapoint used for testing the LGCNN approach.
 Mean $\pm$ standard deviation reported.
-Model Data Case Huber L MAE MSE PAT SSIM
-UNet3dp randomK3 train $0.0025\pm 0.0024$ $4.69\pm 0.17$ $0.0155\pm 0.0071$ $0.0063\pm 0.0062$ $1.92\pm 1.16$ $0.981\pm 0.012$
-val $0.0219\pm 0.0008$ $4.89\pm 0.07$ $0.1114\pm 0.0023$ $0.0474\pm 0.0019$ $34.76\pm 0.99$ $0.696\pm 0.006$
-test $0.0176\pm 0.0006$ $4.85\pm 0.09$ $0.1043\pm 0.0018$ $0.0368\pm 0.0015$ $34.76\pm 0.76$ $0.703\pm 0.006$
-UNet101dp rK101 train $0.0011\pm 0.0002$ $4.37\pm 0.20$ $0.0182\pm 0.0020$ $0.0023\pm 0.0004$ $2.45\pm 0.27$ $0.995\pm 0.002$
-val $0.0055\pm 0.0003$ $4.36\pm 0.16$ $0.0454\pm 0.0019$ $0.0114\pm 0.0006$ $12.96\pm 0.55$ $0.984\pm 0.002$
-test $0.0052\pm 0.0003$ $4.35\pm 0.23$ $0.0441\pm 0.0019$ $0.0110\pm 0.0006$ $12.49\pm 0.48$ $0.985\pm 0.002$
-* LGCNN-test $0.0049\pm 0.0002$ $4.30\pm 0.16$ $0.0470\pm 0.0010$ $0.0102\pm 0.0004$ $13.51\pm 0.59$ $0.983\pm 0.002$
-* scaling $0.0017\pm 0.0001$ $4.48\pm 0.15$ $0.0208\pm 0.0014$ $0.0035\pm 0.0002$ $4.38\pm 0.17$ $0.995\pm 0.001$
-2$\times$2-DDUNet101dp rK101 train $0.0014\pm 0.0003$ $4.11\pm 0.25$ $0.0203\pm 0.0026$ $0.00300\pm 0.0007$ $3.24\pm 0.68$ $0.995\pm 0.001$
-val $0.0079\pm 0.0002$ $4.20\pm 0.25$ $0.0564\pm 0.0008$ $0.01648\pm 0.0005$ $17.32\pm 0.22$ $0.981\pm 0.002$
-test $0.0075\pm 0.0001$ $4.05\pm 0.22$ $0.0552\pm 0.0008$ $0.01580\pm 0.0002$ $16.94\pm 0.36$ $0.982\pm 0.001$
-* LGCNN-test $0.0057\pm 0.0003$ $4.00\pm 0.20$ $0.0526\pm 0.0015$ $0.01171\pm 0.0006$ $16.44\pm 0.63$ $0.981\pm 0.002$
-* scaling $0.0025\pm 0.0001$ $4.04\pm 0.20$ $0.0251\pm 0.0007$ $0.00514\pm 0.0002$ $6.39\pm 0.17$ $0.994\pm 0.001$
-LGCNN randomK3 Step 3 train $0.0001\pm 0.0001$ $2.64\pm 0.18$ $0.0064\pm 0.0008$ $0.0003\pm 0.0001$ $0.29\pm 0.08$ $0.996\pm 0.001$
-val $0.0032\pm 0.0001$ $2.49\pm 0.29$ $0.0413\pm 0.0008$ $0.0065\pm 0.0003$ $10.88\pm 0.35$ $0.912\pm 0.003$
-test $0.0025\pm 0.0000$ $2.74\pm 0.33$ $0.0382\pm 0.0006$ $0.0049\pm 0.0001$ $9.44\pm 0.35$ $0.918\pm 0.003$
-scaling $0.0008\pm 0.0000$ $3.03\pm 0.31$ $0.0179\pm 0.0006$ $0.0016\pm 0.0001$ $2.54\pm 0.14$ $0.946\pm 0.005$
-Table 11: Statistics for predicting $\vec{v}$ with the randomK dataset. Errors in [m/y], MSE in [m2/y2], SSIM unitless. Mean $\pm$ standard deviation reported.
-Model Data Output Case Huber L MAE MSE SSIM
-LGCNN randomK3 $v_{x}$ train 14.11 $\pm$ 11.52 96.63 $\pm$ 23.29 14.60 $\pm$ 11.53 380.19 $\pm$ 529.31 0.997 $\pm$ 0.002
-$v_{y}$ train 14.30 $\pm$ 7.10 135.26 $\pm$ 20.36 14.80 $\pm$ 7.11 361.61 $\pm$ 296.21 0.991 $\pm$ 0.006
-$v_{x}$ val 28.23 $\pm$ 7.76 445.13 $\pm$ 35.50 28.72 $\pm$ 7.77 1601.10 $\pm$ 587.80 0.989 $\pm$ 0.003
-$v_{y}$ val 27.33 $\pm$ 2.25 343.35 $\pm$ 26.80 27.82 $\pm$ 2.25 1587.57 $\pm$ 252.10 0.982 $\pm$ 0.002
-$v_{x}$ test 26.94 $\pm$ 5.99 216.74 $\pm$ 18.92 27.44 $\pm$ 5.99 1368.52 $\pm$ 431.48 0.990 $\pm$ 0.002
-$v_{y}$ test 30.38 $\pm$ 6.18 249.11 $\pm$ 11.37 30.88 $\pm$ 6.18 1848.47 $\pm$ 637.10 0.982 $\pm$ 0.005
-$v_{x}$ scaling 27.89 $\pm$ 6.53 286.08 $\pm$ 22.96 28.38 $\pm$ 6.53 1486.23 $\pm$ 451.52 0.990 $\pm$ 0.002
-$v_{y}$ scaling 27.60 $\pm$ 2.43 328.80 $\pm$ 24.95 28.09 $\pm$ 2.43 1578.87 $\pm$ 258.66 0.979 $\pm$ 0.003
-LGCNN realK $v_{x}$ train 29.66 $\pm$ 2.29 592.17 $\pm$ 37.66 30.16 $\pm$ 2.29 2326.33 $\pm$ 327.11 0.988 $\pm$ 0.005
-$v_{y}$ train 28.93 $\pm$ 5.47 397.09 $\pm$ 72.78 29.43 $\pm$ 5.47 1778.28 $\pm$ 551.97 0.997 $\pm$ 0.009
-$v_{x}$ val 22.17 $\pm$ 2.24 170.47 $\pm$ 9.75 22.66 $\pm$ 2.24 909.39 $\pm$ 139.94 0.988 $\pm$ 0.002
-$v_{y}$ val 20.20 $\pm$ 4.07 113.33 $\pm$ 12.37 20.70 $\pm$ 4.08 645.11 $\pm$ 221.50 0.997 $\pm$ 0.001
-$v_{x}$ scaling 79.33 $\pm$ 21.80 532.35 $\pm$ 140.50 79.83 $\pm$ 21.80 9404.62 $\pm$ 4048.09 0.953 $\pm$ 0.014
-$v_{y}$ scaling 53.39 $\pm$ 12.64 602.41 $\pm$ 128.79 53.89 $\pm$ 12.64 4621.25 $\pm$ 1755.50 0.980 $\pm$ 0.006
+Model Data Case Huber L MAE MSE
+UNet 101dp train $0.0011\pm 0.0002$ $4.37\pm 0.20$ $0182\pm 0.0020 0.0$$0023\pm 0.0004$
+val $0.0055\pm 0.0003$ $4.36\pm 0.16$ $0454\pm 0.0019 0.0$$0114\pm 0.0006$
+test $0.0052\pm 0.0003$ $4.35\pm 0.23$ $0441\pm 0.0019 0.0$$0110\pm 0.0006$
+* LGCNN-test $0.0049\pm 0.0002$ $4.30\pm 0.16$ $0470\pm 0.0010 0.0$$0102\pm 0.0004$
+* scaling $0.0017\pm 0.0001$ $4.48\pm 0.15$ $0208\pm 0.0014 0.0$$0035\pm 0.0002$
+DDUNet 101dp train $0.0014\pm 0.0003$ $4.11\pm 0.25$ $0203\pm 0.0026 0.0$$00300\pm 0.0007$
+val $0.0079\pm 0.0002$ $4.20\pm 0.25$ $0564\pm 0.0008 0.0$$01648\pm 0.0005$
+test $0.0075\pm 0.0001$ $4.05\pm 0.22$ $0552\pm 0.0008 0.0$$01580\pm 0.0002$
+* LGCNN-test $0.0057\pm 0.0003$ $4.00\pm 0.20$ $0526\pm 0.0015 0.0$$01171\pm 0.0006$
+* scaling $0.0025\pm 0.0001$ $4.04\pm 0.20$ $0251\pm 0.0007 0.0$$00514\pm 0.0002$
+Vanilla UNet on 3dp: Metrics of training and ablation study
+We present metrics for predicting $T$ directly from the inputs pkipki with UNet3dp in Table˜8, evaluated on training, validation, test and scaling datapoints.
+To assess the model’s sensitivity to random initialization, the training of the same architecture was repeated five times.
+Other models were not trained again to reduce computational effort and because the results of all vanilla approaches indicate that the performance of UNet101dp, UNet3dp, and DDUNet101dp can be viewed as a representative of the general performance of purely data-driven methods.
+Table 8: Statistical error metrics for predicting with different models and datasets. Errors in [°C] or [m/y], MoC in [%], and SSIM unitless. Mean $\pm$ standard deviation reported.
+Architecture Data Split Scaling Step MAE RMSE Huber SSIM MoC
+UNeta synth.: 3dp train No end-to-end 0.0155 $\pm$ 0.0064 0.0732 $\pm$ 0.0310 0.0025 $\pm$ 0.0021 0.9645 $\pm$ 0.0157 1.6770 $\pm$ 1.9093
+UNeta synth.: 3dp val No end-to-end 0.1114 $\pm$ 0.0021 0.2177 $\pm$ 0.0039 0.0219 $\pm$ 0.0007 0.6251 $\pm$ 0.0067 6.9180 $\pm$ 0.7236
+UNeta synth.: 3dp test No end-to-end 0.1043 $\pm$ 0.0016 0.1917 $\pm$ 0.0035 0.0176 $\pm$ 0.0006 0.6239 $\pm$ 0.0048 7.7775 $\pm$ 1.5033
+LGCNNb synth.: 3dp train No end-to-end 0.0720 $\pm$ 0.0012 0.1736 $\pm$ 0.0019 0.0137 $\pm$ 0.0003 0.7993 $\pm$ 0.0044 0.0933 $\pm$ 0.0430
+LGCNNb synth.: 3dp val No end-to-end 0.1026 $\pm$ 0.0014 0.2095 $\pm$ 0.0032 0.0204 $\pm$ 0.0006 0.6812 $\pm$ 0.0061 1.6610 $\pm$ 0.0797
+LGCNNb synth.: 3dp test No end-to-end 0.0965 $\pm$ 0.0013 0.1866 $\pm$ 0.0030 0.0167 $\pm$ 0.0005 0.6724 $\pm$ 0.0059 0.1443 $\pm$ 0.0388
+LGCNNb synth.: 3dp test Yes end-to-end 0.0430 $\pm$ 0.0007 0.1249 $\pm$ 0.0016 0.0072 $\pm$ 0.0002 0.7908 $\pm$ 0.0151 0.0951 $\pm$ 0.0317
+UNet synth.: 3dp train No 3 0.0354 $\pm$ 0.0007 0.0770 $\pm$ 0.0024 0.0029 $\pm$ 0.0002 0.9084 $\pm$ 0.0032 0.1516 $\pm$ 0.0203
+UNet synth.: 3dp val No 3 0.0491 $\pm$ 0.0005 0.0916 $\pm$ 0.0008 0.0042 $\pm$ 0.0001 0.8251 $\pm$ 0.0033 0.1970 $\pm$ 0.0381
+UNet synth.: 3dp test No 3 0.0452 $\pm$ 0.0005 0.0829 $\pm$ 0.0011 0.0034 $\pm$ 0.0001 0.8381 $\pm$ 0.0040 0.1938 $\pm$ 0.0907
+UNet synth.: 3dp test Yes 3 0.0179 $\pm$ 0.0005 0.0399 $\pm$ 0.0007 0.0008 $\pm$ 0.0000 0.8809 $\pm$ 0.0138 0.1148 $\pm$ 0.0497
+UNet synth.: 3dp train No 1 (vx) 14.6004 $\pm$ 10.3108 16.7442 $\pm$ 9.9912 14.1096 $\pm$ 10.3058 0.9884 $\pm$ 0.0020 -
+UNet synth.: 3dp train No 1 (vy) 14.7965 $\pm$ 6.3568 17.9164 $\pm$ 6.3728 14.3037 $\pm$ 6.3526 0.9868 $\pm$ 0.0051 -
+UNet synth.: 3dp val No 1 (vx) 28.7237 $\pm$ 6.9461 39.5616 $\pm$ 5.9981 28.2283 $\pm$ 6.9444 0.9734 $\pm$ 0.0025 -
+UNet synth.: 3dp val No 1 (vy) 27.8241 $\pm$ 2.0130 39.7497 $\pm$ 2.7446 27.3289 $\pm$ 2.0125 0.9743 $\pm$ 0.0020 -
+UNet synth.: 3dp test No 1 (vx) 27.4355 $\pm$ 5.3603 36.6786 $\pm$ 4.8168 26.9402 $\pm$ 5.3588 0.9727 $\pm$ 0.0025 -
+UNet synth.: 3dp test No 1 (vy) 30.8755 $\pm$ 5.5264 42.5114 $\pm$ 6.4231 30.3799 $\pm$ 5.5256 0.9734 $\pm$ 0.0042 -
+UNet synth.: 3dp test Yes 1 (vx) 28.3820 $\pm$ 5.8382 38.2430 $\pm$ 4.8683 27.8864 $\pm$ 5.8367 0.9770 $\pm$ 0.0017 -
+UNet synth.: 3dp test Yes 1 (vy) 28.0927 $\pm$ 2.1725 39.6320 $\pm$ 2.8592 27.5972 $\pm$ 2.1719 0.9717 $\pm$ 0.0025 -
+LGCNNb real: 4dp train No end-to-end 0.0550 $\pm$ 0.0025 0.0960 $\pm$ 0.0014 0.0046 $\pm$ 0.0001 0.6744 $\pm$ 0.0249 1.8951 $\pm$ 0.6214
+LGCNNb real: 4dp val No end-to-end 0.0827 $\pm$ 0.0029 0.1553 $\pm$ 0.0031 0.0120 $\pm$ 0.0005 0.6337 $\pm$ 0.0195 2.0679 $\pm$ 0.4008
+LGCNNb real: 4dp test Yes end-to-end 0.0486 $\pm$ 0.0075 0.0714 $\pm$ 0.0056 0.0026 $\pm$ 0.0004 0.6257 $\pm$ 0.0396 2.7049 $\pm$ 0.6239
+UNet real: 4dp train No 3 0.0213 $\pm$ 0.0051 0.0295 $\pm$ 0.0061 0.0005 $\pm$ 0.0002 0.8895 $\pm$ 0.0208 1.9966 $\pm$ 0.2618
+UNet real: 4dp val No 3 0.0239 $\pm$ 0.0023 0.0395 $\pm$ 0.0034 0.0008 $\pm$ 0.0001 0.8982 $\pm$ 0.0169 2.9922 $\pm$ 0.3038
+UNet real: 4dp test Yes 3 0.0218 $\pm$ 0.0042 0.0326 $\pm$ 0.0044 0.0005 $\pm$ 0.0001 0.8734 $\pm$ 0.0190 1.6682 $\pm$ 0.2054
+UNet real: 4dp train No 1 (vx) 22.0019 $\pm$ 5.7393 27.7180 $\pm$ 6.8232 21.5061 $\pm$ 5.7380 0.9731 $\pm$ 0.0280 -
+UNet real: 4dp train No 1 (vy) 13.4570 $\pm$ 9.2203 16.8320 $\pm$ 10.8123 12.9699 $\pm$ 9.2128 0.9607 $\pm$ 0.0629 -
+UNet real: 4dp val No 1 (vx) 24.8187 $\pm$ 4.0315 29.6583 $\pm$ 4.7384 24.3217 $\pm$ 4.0310 0.9641 $\pm$ 0.0220 -
+UNet real: 4dp val No 1 (vy) 12.9956 $\pm$ 9.9800 15.4509 $\pm$ 10.3446 12.5082 $\pm$ 9.9727 0.9485 $\pm$ 0.0748 -
+UNet real: 4dp test Yes 1 (vx) 122.7812 $\pm$ 32.6445 135.7671 $\pm$ 36.8100 122.2815 $\pm$ 32.6443 0.9128 $\pm$ 0.0511 -
+UNet real: 4dp test Yes 1 (vy) 25.5754 $\pm$ 6.2873 34.6947 $\pm$ 8.2358 25.0801 $\pm$ 6.2858 0.9445 $\pm$ 0.0834 -
+a newly trained with other validation loss for comparability in this table
+b with fixed first step predictions
 LGCNN: Metrics of training and ablation study
-The results of the LGCNN and DDU-Net, evaluated on both synthetic and realistic permeability fields, for the training, validation, testing, and scaling datasets are shown in Table 12 (for Step 1 - predict $\vec{v}_{\text{pred}}$) and Table 13 (for Step 3 - predict $T$).
-The column "data" refers to the dataset that a model was trained on and applied to, i.e., "randomK" stands for the baseline dataset 3dp, "randomK101" for 101dp.
-An asterisk indicates that the model was evaluated on a dataset different from the one it was trained on.
-Mostly relevant for the vanilla approaches that are trained on 101dp and applied to scaling and test of 3dp.
-Table 12: Performance metrics for for predicting $\vec{v}$ with different models and datasets. Errors in [m/y], MSE in [m2/y2], SSIM unitless.
-Model Data Output Case Huber L MAE MSE SSIM
-1st Step
-LGCNN randomK $v_{x}$ train 9.4732 132.6672 9.9620 171.0032 0.9972
-$v_{y}$ train 11.4005 223.4601 11.8902 275.6190 0.9937
-$v_{x}$ val 22.2241 343.9907 22.7179 1102.3721 0.9905
-$v_{y}$ val 26.6078 274.8036 27.1026 1524.5099 0.9841
-$v_{x}$ test 21.8237 190.8046 22.3178 972.5668 0.9911
-$v_{y}$ test 32.2488 256.2519 32.7444 2031.3357 0.9812
-* $v_{x}$ scaling 24.4314 294.0457 24.9261 1204.1154 0.9911
-* $v_{y}$ scaling 25.7847 367.6891 26.2795 1463.8218 0.9820
-Experiment: trained on full image
-LGCNN randomK $v_{x}$ train 3.1983 39.9380 3.6714 21.5058 0.9993
-$v_{y}$ train 3.0496 43.9018 3.5177 21.4905 0.9988
-$v_{x}$ val 25.2418 417.1447 25.7367 1421.6547 0.9896
-$v_{y}$ val 29.6294 276.8029 30.1247 1827.3134 0.9824
-$v_{x}$ test 23.3646 202.6397 23.8584 1144.4479 0.9913
-$v_{y}$ test 35.4289 332.9557 35.9249 2490.4104 0.9789
-Replace with UNet-101dp, i.e., $1\times 1$ subdomain-DDUNet, trained on 101 datapoints
-DDUNet randomK101 $v_{x}$ train 3.5468 678.7761 4.0134 86.6731 0.9767
-$v_{y}$ train 4.1030 690.1506 4.5816 100.2767 0.9934
-$v_{x}$ val 3.8368 647.1833 4.3058 90.3292 0.9739
-$v_{y}$ val 4.3649 685.6656 4.8433 102.8601 0.9931
-$v_{x}$ test 3.6643 584.8119 4.1325 83.4062 0.9731
-$v_{y}$ test 4.1355 547.8865 4.6131 87.7819 0.9924
-* $v_{x}$ LGCNN-test 12.2161 1047.8464 12.7036 896.2758 0.9016
-* $v_{y}$ LGCNN-test 13.9033 2594.9504 14.3878 2317.6863 0.9903
-* $v_{x}$ scaling 4.7412 659.8439 5.2152 79.2508 0.9699
-* $v_{y}$ scaling 5.2686 456.1904 5.7492 83.9735 0.9948
-Replace with 2x2-DDUNet-101dp, i.e.,2×$\times 2$ 2 subdomains
-DDUNet randomK101 $v_{x}$ train 5.3928 422.4526 5.8815 57.6045 0.9580
-$v_{y}$ train 4.0476 696.2919 4.5116 58.8457 0.9972
-$v_{x}$ val 5.7663 471.6309 6.2545 76.4372 0.9552
-$v_{y}$ val 4.5085 644.1844 4.9751 71.5679 0.9973
-$v_{x}$ test 5.5847 319.9844 6.0735 59.2501 0.9526
-$v_{y}$ test 4.2292 535.0656 4.6948 54.5020 0.9972
-* $v_{x}$ LGCNN-test 13.1829 2453.5881 13.6735 1367.9061 0.9325
-* $v_{y}$ LGCNN-test 14.6481 2029.0269 15.1332 2221.4226 0.9919
-* $v_{x}$ scaling 7.6311 421.4186 8.1184 125.9511 0.9314
-* $v_{y}$ scaling 6.9406 400.2490 7.4200 135.7929 0.9955
-Domain Transfer, i.e., to more realistic data
-LGCNN realK $v_{x}$ train 13.6890 122.7981 14.1819 340.6713 0.9973
-$v_{y}$ train 9.0680 71.2576 9.5619 126.7222 0.9991
-$v_{x}$ val 14.9187 106.8607 15.4095 380.3762 0.9939
-$v_{y}$ val 10.1675 74.5570 10.6605 148.2475 0.9993
-* $v_{x}$ scaling 109.5079 459.1620 110.0078 13570.0000 0.9462
-* $v_{y}$ scaling 17.1118 240.9406 17.6051 616.5186 0.9965
-Table 13: Performance metrics for predicting $T$ with different models and datasets. Errors in [°C], MSE in [°C2], PATs in [%], SSIM unitless.
-Model Data Case Huber L MAE MSE PAT SSIM
-3rd Step: trained and applied to $v_{sim}$
-LGCNN randomK $v_{sim}$ train 1.87e-5 2.2536 0.0014 3.95e-5 0.03 0.9997
-val 0.0027 2.8857 0.0369 0.0054 8.61 0.9283
-test 0.0021 2.8990 0.0347 0.0041 7.54 0.9304
-* scaling 0.0007 3.0250 0.0168 0.0014 2.05 0.9510
-Full Pipeline, i.e., Step 3 trained on $v_{sim}$, but applied to $v_{pred}$
-LGCNN randomK $v_{pred}$ train 0.0121 5.1006 0.0642 0.0272 18.17 0.8700
-val 0.0188 4.2264 0.0967 0.0411 28.96 0.7625
-test 0.0147 4.2120 0.0905 0.0307 28.92 0.7637
-* scaling 0.0065 4.9366 0.0413 0.0141 10.87 0.8654
-Experiments on 3rd Step
-Trained in sequence, i.e., trained on $v_{pred}$
-LGCNN randomK $v_{pred}$ train 0.0032 4.0750 0.0198 0.0068 3.73 0.9794
-val 0.0195 3.8178 0.1003 0.0423 30.46 0.7442
-test 0.0139 4.0804 0.0901 0.0289 29.44 0.7553
-Trained with zero-padding
-LGCNN randomK $v_{sim}$ train 0.0002 1.8465 0.0140 0.0004 0.33 0.9518
-val 0.0032 2.9051 0.0465 0.0064 11.31 0.8686
-test 0.0024 2.3391 0.0418 0.0048 9.19 0.8838
-Trained on full image
-LGCNN randomK $v_{sim}$ train 0.0002 3.5648 0.0022 0.0004 0.09 0.9992
-val 0.0027 3.5088 0.0375 0.0055 8.81 0.9217
-test 0.0021 3.6420 0.0351 0.0043 7.41 0.9245
-Replace with UNet-101dp: Step 3 trained and applied on $v_{sim}$
-DDUNet rK101 train 0.0019 3.8491 0.0283 0.0038 5.35 0.9911
-val 0.0027 3.9381 0.0350 0.0055 8.89 0.9895
-test 0.0025 3.7999 0.0343 0.0052 8.50 0.9903
-* LGCNN-test 0.0032 4.4658 0.0416 0.0067 11.69 0.9835
-* scaling 0.0010 4.3489 0.0176 0.0022 3.07 0.9954
-Replace with 2x2-DDUNet-101dp: Step 3 trained and applied on $v_{sim}$
-DDUNet rK101 train 0.0015 4.0624 0.0200 0.0032 3.44 0.9952
-val 0.0040 4.1863 0.0410 0.0083 12.54 0.9904
-test 0.0039 4.0807 0.0411 0.0081 12.67 0.9910
-* LGCNN-test 0.0029 4.2772 0.0376 0.0058 10.96 0.9915
-* scaling 0.0012 3.9946 0.0192 0.0025 4.73 0.9968
-Domain Transfer, i.e. to more realistic data
-3rd Step
-LGCNN realK $v_{sim}$ train 0.0002 0.7704 0.0139 0.0005 0.43 2.9107
-val 0.0005 0.8222 0.0175 0.0010 2.18 0.9672
-* scaling 0.0004 0.8052 0.0189 0.0008 0.92 0.9497
-Full Pipeline
-LGCNN realK $v_{pred}$ train 0.0049 2.5437 0.0534 0.0100 17.58 2.4923
-val 0.0137 2.3194 0.0841 0.0275 27.79 0.7510
-* scaling 0.0022 2.0511 0.0394 0.0044 10.02 0.8708
+The results of all steps/end-to-end, evaluated on both synthetic and realistic permeability fields, for the training, validation, testing, and scaling datasets are shown in Table˜9.
+Table 9: Performance metrics for predicting with different models and datasets. Errors in [°C] or [m/y], MoC in [%], and SSIM unitless.
+Architecture Data Split Scaling Step MAE RMSE Huber SSIM MoC
+UNet synthetic: 3dp train No end-to-end 0.1181 0.5537 0.0558 0.8454 2.29
+UNet synthetic: 3dp val No end-to-end 0.1930 0.5795 0.0742 0.6515 11.32
+UNet synthetic: 3dp test No end-to-end 0.1908 0.6190 0.0736 0.6671 10.85
+DilatedCNN synthetic: 3dp train No end-to-end 0.1180 0.2494 0.0270 0.5483 11.28
+DilatedCNN synthetic: 3dp val No end-to-end 0.1273 0.2513 0.0287 0.5183 10.67
+DilatedCNN synthetic: 3dp test No end-to-end 0.1200 0.2236 0.0237 0.5004 10.95
+PINN (overfit) synthetic: 1dp train No 3 0.7532 0.8593 0.3454 0.1793 0
+LGCNN synthetic: 3dp train No end-to-end 0.0657 0.1640 0.0121 0.8209 0.08
+LGCNN synthetic: 3dp val No end-to-end 0.0965 0.1996 0.0184 0.6933 1.54
+LGCNN synthetic: 3dp test No end-to-end 0.0916 0.1738 0.0146 0.6841 0.09
+LGCNN synthetic: 3dp test Yes end-to-end 0.0413 0.1189 0.0065 0.7911 0.06
+UNet synthetic: 3dp train No 1 (vx) 9.9620 13.0768 9.4732 0.9868 -
+UNet synthetic: 3dp val No 1 (vx) 22.7179 33.2020 22.2241 0.9748 -
+UNet synthetic: 3dp test No 1 (vx) 22.3178 31.1860 21.8237 0.9739 -
+UNet synthetic: 3dp test Yes 1 (vx) 24.9261 34.7004 24.4314 0.9784 -
+UNet synthetic: 3dp train No 1 (vy) 11.8902 16.6018 11.4005 0.9890 -
+UNet synthetic: 3dp val No 1 (vy) 27.1026 39.0450 26.6078 0.9770 -
+UNet synthetic: 3dp test No 1 (vy) 32.7444 45.0703 32.2488 0.9733 -
+UNet synthetic: 3dp test Yes 1 (vy) 26.2795 38.2599 25.7847 0.9761 -
+UNet synthetic: 3dp train No 3 0.0282 0.0632 0.0020 0.9376 0.11
+UNet synthetic: 3dp val No 3 0.0452 0.0869 0.0038 0.8454 0.11
+UNet synthetic: 3dp test No 3 0.0417 0.0762 0.0029 0.8540 0.07
+UNet synthetic: 3dp test Yes 3 0.0168 0.0373 0.0007 0.8895 0.08
+UNet, in sequence synthetic: 3dp train No 3 0.0402 0.1039 0.0052 0.9019 0.08
+UNet, in sequence synthetic: 3dp val No 3 0.1013 0.2025 0.0191 0.6682 1.15
+UNet, in sequence synthetic: 3dp test No 3 0.0919 0.1695 0.0139 0.6714 1.30
+UNet, with zero-padding synthetic: 3dp train No 3 0.0440 0.0788 0.0031 0.8104 0.44
+UNet, with zero-padding synthetic: 3dp val No 3 0.0548 0.0908 0.0041 0.7546 0.70
+UNet, with zero-padding synthetic: 3dp test No 3 0.0487 0.0795 0.0031 0.7784 0.57
+UNet, no partitioning synthetic: 3dp train No 1 (vx) 3.6714 4.6374 3.1983 0.9959 -
+UNet, no partitioning synthetic: 3dp val No 1 (vx) 25.7367 37.7048 25.2418 0.9764 -
+UNet, no partitioning synthetic: 3dp test No 1 (vx) 23.8584 33.8297 23.3646 0.9765 -
+UNet, no partitioning synthetic: 3dp train No 1 (vy) 3.5177 4.6358 3.0496 0.9971 -
+UNet, no partitioning synthetic: 3dp val No 1 (vy) 30.1247 42.7471 29.6294 0.9762 -
+UNet, no partitioning synthetic: 3dp test No 1 (vy) 35.9249 49.9040 35.4289 0.9721 -
+UNet, no partitioning synthetic: 3dp train No 3 0.0323 0.0736 0.0027 0.9245 0.14
+UNet, no partitioning synthetic: 3dp val No 3 0.0449 0.0861 0.0037 0.8333 0.12
+UNet, no partitioning synthetic: 3dp test No 3 0.0429 0.0793 0.0031 0.8349 0.29
+UNet real: 4dp train No 1 (vx) 14.1819 18.4573 13.6890 0.9915 -
+UNet real: 4dp val No 1 (vx) 15.4095 19.5032 14.9187 0.9869 -
+UNet real: 4dp test Yes 1 (vx) 110.0078 116.4903 109.5079 0.9436 -
+UNet real: 4dp train No 1 (vy) 9.5619 11.2571 9.0680 0.9945 -
+UNet real: 4dp val No 1 (vy) 10.6605 12.1757 10.1675 0.9932 -
+UNet real: 4dp test Yes 1 (vy) 17.6051 24.8298 17.1118 0.9905 -
+UNet real: 4dp train No 3 0.0139 0.0212 0.0002 0.9361 1.88
+UNet real: 4dp val No 3 0.0175 0.0319 0.0005 0.9405 3.30
+UNet real: 4dp test Yes 3 0.0189 0.0287 0.0004 0.8942 1.77
+LGCNN real: 4dp train No end-to-end 0.0534 0.0997 0.0049 0.6900 2.76
+LGCNN real: 4dp val No end-to-end 0.0841 0.1659 0.0137 0.6511 1.65
+LGCNN real: 4dp test Yes end-to-end 0.0394 0.0665 0.0022 0.6755 3.50
 LGCNN+random $k$: Performance of Step 1
-The model generally obtains good results in Figure 3, even for cells that are far away from injection points.
+The model generally obtains good results in Figure˜3, even for cells that are far away from injection points.
 Figure 6: Prediction of $\vec{v}$ with LGCNN. Left: Test datapoint with labels $v_{x}$ and $v_{y}$. Middle: Step 1 prediction for $v_{x}$, $v_{y}$. Right: Error of the $v_{x}$, $v_{y}$ prediction.
@@
 Nonetheless, we retain the implicit scheme for its superior accuracy and stability.
-Quantitative and qualitative results for the predictions are shown in Figure 7 and Table 14.
-Table 14: Experiment on 3rd step: Test metrics for predicting $T$ with different input combinations. Errors in [°C], MSE in [°C2], PAT in [%], SSIM unitless.
-Inputs Huber L MAE MSE PAT SSIM
-$ikv_{x}v_{y}$ 0.0070 2.2990 0.0712 0.0139 25.55 0.7662
-$ikv_{x}v_{y}s$ 0.0041 1.8674 0.0545 0.0083 20.56 0.8368
-$ikv_{x}v_{y}s_{o}$ 0.0057 2.6623 0.0598 0.0114 20.64 0.8423
-$ikv_{x}v_{y}ss_{o}$ (not faded) 0.0072 2.3039 0.0744 0.0144 21.40 0.7837
-$ikv_{x}v_{y}ss_{o}^{\mathrm{a}}$ 0.0031 1.8364 0.0442 0.0062 12.97 0.8828
-$v_{x}v_{y}ss_{o}$ 0.0066 2.1301 0.0647 0.0132 20.64 0.8681
-$kv_{x}v_{y}ss_{o}$ 0.0049 2.0925 0.0587 0.0097 20.32 0.8732
-explicit RK, order 4 0.0038 3.2636 0.0486 0.0076 15.80 0.8871
-explicit RK, order 2 0.0036 2.2034 0.0463 0.0072 13.75 0.8830
+Quantitative and qualitative results for the predictions are shown in Figure˜7 and Table˜10.
+Table 10: Experiment on 3rd step: Test metrics for predicting $T$ with different input combinations. Errors in [°C], MSE in [°C2], and SSIM unitless.
+Inputs Huber L MAE MSE SSIM
+$ikv_{x}v_{y}$ 0.0070 2.2990 0.0712 0.0139 0.7662
+$ikv_{x}v_{y}s$ 0.0041 1.8674 0.0545 0.0083 0.8368
+$ikv_{x}v_{y}s_{o}$ 0.0057 2.6623 0.0598 0.0114 0.8423
+$ikv_{x}v_{y}ss_{o}$ (not faded) 0.0072 2.3039 0.0744 0.0144 0.7837
+$ikv_{x}v_{y}ss_{o}^{\mathrm{a}}$ 0.0031 1.8364 0.0442 0.0062 0.8828
+$v_{x}v_{y}ss_{o}$ 0.0066 2.1301 0.0647 0.0132 0.8681
+$kv_{x}v_{y}ss_{o}$ 0.0049 2.0925 0.0587 0.0097 0.8732
+explicit RK, order 4 0.0038 3.2636 0.0486 0.0076 0.8871
+explicit RK, order 2 0.0036 2.2034 0.0463 0.0072 0.8830
 a new run to be comparable to the others in this experiment: trained
-with Huber validation loss, hence the results differ slightly wrt. to Table 3.
+with Huber validation loss, hence the results differ slightly wrt. to Table˜1.
 Figure 7: 1st Column: Label, input $k$, 2nd-5th: Without $(s,s_{o})$, $(s_{o})$, $(s)$, not-faded streamlines $(s,s_{o})$, 6th: include all inputs, 7th-8th: Without $(i,k)$, $(i)$.
 Absolute errors capped at 1°C for better visualizations.
-Maximum errors are listed in Table 14.
+Maximum errors are listed in Table˜10.
 LGCNN+realistic $k$: Performance of 3rd step and full pipeline
-The qualitative performance is observable in Figure 8, where we see coherent streamlines and plume structures for both the isolated 3rd step and the full pipeline.
+The qualitative performance is observable in Figure˜8, where we see coherent streamlines and plume structures for both the isolated 3rd step and the full pipeline.
 Figure 8: 1st Column: Label, input $k$. 2nd Column: 3rd step prediction of $T$ and error with respect to the label. 3rd Column: Predicted $T$ and error of the full pipeline.
 Training and inference times
-Table 15 summarizes the training and inference times, number of epochs, and dataset splits (train:val:test) for each of the three steps in our pipeline, both for the LGCNN trained on partitioned and full datasets.
+Table˜11 summarizes the training and inference times, number of epochs, and dataset splits (train:val:test) for each of the three steps in our pipeline, both for the LGCNN trained on partitioned and full datasets.
 Although both approaches exhibit similar inference times, they show significant differences in training time, with the partitioned approach yielding better results.
-Table 15: Training measurements on the randomK dataset.
+Table 11: Training measurements on 3dp dataset.
 Data Points Epochsa Training Timea Inference Time
@@
 aEarly stopping: measurements until best validation loss.
-In Table 16, the number of epochs and the total training time for the data-driven approaches are shown.
+In Table˜12, the number of epochs and the total training time for the data-driven approaches are shown.
 For Step 1, both the UNet and DDUNet need many epochs and comparable training time to converge; however, for the third step and the full pipeline, the DDUNet significantly reduces both the number of epochs and the total training time required to reach convergence.
@@
 UNet101dp 738 5.497 hours
-2$\times$2 DDUNet101dp 735 4.787 hours
+$\times2$ DDUNet101dp 735 4.787 hours
 3rd Step
 UNet101dp 726 8.343 hours
-2$\times$2 DDUNet101dp 303 3.508 hours
+$\times2$ DDUNet101dp 303 3.508 hours
 Full Pipeline
 UNet101dp 267 3.680 hours
-2$\times$2 DDUNet101dp 97 1.222 hours
+$\times2$ DDUNet101dp 97 1.222 hours
 aEarly stopping: measurements until best validation loss.
 A.4 Hardware Specifications
-The 2$\times$2 DDU-Net101dp and UNet101dp models, trained on the large data-driven dataset of 101 samples, were trained and evaluated on the DelftBlue cluster [DHPC2024], using NVIDIA V100 GPUs with 32 GB memory.
+The vanilla models that were trained on the large data-driven dataset of 101 samples were trained and evaluated on a server using NVIDIA V100 GPUs with 32 GB memory.
 All training was conducted using PyTorch 2.1.0 with CUDA 11.6 acceleration.
-Training and evaluation of the LGCNN model were performed on a single NVIDIA A100-SXM4 GPU. Data generation was carried out on a dual-socket system equipped with AMD EPYC 9274F CPUs.
+Training and evaluation of all other models (dilated UNet, LGCNN single steps, full pipeline and ablation studies, FNOs, PINNs) were performed on a single NVIDIA A100-SXM4 GPU.
+Data generation was carried out on a dual-socket system equipped with AMD EPYC 9274F CPUs.
 A.5 Glossary
-A list of the most relevant physical properties used in our paper is provided in Table 17.
+A list of the most relevant physical properties used in our paper is provided in Table˜13.
 Table 13: Glossary of Abbreviations.
@@
 $X_{pred}$ predicted property X
+$X_{sim}$ simulated property X
 $i$ positions of heat pumps
@@
 $\vec{s}$ both streamline fields
-$s$ central streamlines after all $i$
-$s_{o}$ streamlines with transversal offset to $i$
+$s$ central streamlines, starting at all $i$
+$s_{o}$ streamlines starting with $\pm$ 10 cells transversal offset compared to $s$
 $T$ temperature